### v3

# The Poisson Follower Model

Abstract
We introduce a stochastic geometry dynamics inspired by opinion dynamics that captures the essence of modern asymmetric social networks with leaders and followers. Opinions are represented by points in Euclidean space, and the leader of an agent is the one with the closest opinion. In this dynamics, each follower updates its opinion by halving the distance to its leader. We demonstrate that this simple dynamics and its iterations exhibit several interesting purely geometric phenomena related to the evolution of leadership and opinion clusters, which resemble those observed in social networks. We also show that when the initial opinions are randomly distributed as a stationary Poisson point process, the likelihood of each of these phenomena can be expressed through an integral geometry formula involving semi-algebraic domains. Furthermore, we establish this property for step 0 and step 1 of the dynamics using percolation techniques. Finally, we analyze numerically the limiting behavior of this follower dynamics. In the Poisson case, the agents fall into two categories: ultimate followers, who continue updating their opinions indefinitely, and ultimate leaders, who adopt a fixed opinion after a finite time. All our findings are supported by spatial discrete event simulations.
1 Introduction
People’s opinions and beliefs are influenced in complex ways by families, friends, colleagues, and media, as well as by politicians and other mega-influencers [1, 10, 11, 32, 21, 24]. In recent decades, attempts have been made to understand aspects of this process using mathematical modeling and computational simulation such as surveys on opinion dynamics [3, 2, 50, 55, 61, 62].
Many models of opinion dynamics are based on the assumption that we are influenced more easily by people with whom we almost agree than by those whose views starkly differ from ours. A similar but more general phenomenon, known as biased assimilation, is recognized among psychologists; referring to our tendency to filter and interpret information in a way that supports our preconceived notions [47]. Models of opinion dynamics built on this assumption are known as bounded confidence models [3, 23, 53]. A popular example comes from Hegselmann and Krause [35, 36], which builds on earlier work by Krause [40, 41]. It has been extensively studied in the literature [48, 49, 14, 63], and will serve as our starting point here. The Hegselmann-Krause model examines the stochastic time evolution of a voter’s opinion in response to the opinion of other like-minded voters.
The original Hegselmann-Krause model is discrete in both time and opinion space with a finite number of agents moving one step at a time. Similar models that are continuous in time [58], opinion space [69], or opinion space and time [19, 20, 30] have been proposed since then.
This paper proposes a mathematical model allowing one to analyze both short and long term opinion dynamics of a countable collection of agents. At any time step of the dynamics, each agent has an $\mathbb{R}^{d}$-valued opinion and can only be influenced by the agent whose opinion is closest to its opinion, which is referred to as its leader. This new model is meant to capture the essence of such social network interactions as Instagram and Twitter, where the asymmetric leader/follower dynamics is central.
The follower/leader connections form a directed graph, where the direction goes from the follower towards the leader. Two agents which follow each other will be called an ultimate leader pair. We think of a group of agents that are part of the same connected component of this follower graph as a party.
If the initial opinions form a random point process, the directed graph is generated over a random point set, and is hence a random graph. In several parts of this work, we will restrict ourselves to the case where the points of the initial configuration are sampled according to a stationary Poisson point process over $\mathbb{R}^{d}$. This choice is dictated by tractability reasons (other initial conditions will be considered in future steps). For instance, in this Poisson case, the follower/leader random graph is called the Poisson Nearest Neighbor graph and parties are called Poisson descending trees. They have been studied extensively [33, 46]. An example of such a tree is depicted in Figure 1 for $d=2$.
Figure 1: Example of a Poisson descending tree
The initial random graph evolves over time based on the prescribed dynamics. More precisely, given an initial point process, at the first time step of the dynamics, each agent’s opinion moves halfway to that of its leader. This gives a second point process on which a new follower/leader directed graph is defined. The dynamics is then applied to this new graph, which defines a third point process, and so on. The Poisson Follower Point Process of order $n$ is the point process obtained at the $n^{th}$ iteration of this dynamics when starting from a Poisson Point Process. At the $n^{th}$ iteration, the agents or points are the vertices of a graph that will be referred to as the Poisson Follower graph of order $n$. As we will see, a given agent may follow different leaders at different times, the structure of the Poisson follower graph changes with $n$.
This dynamics can be seen as a variant of the Hegselmann-Krause dynamics [35]. The main difference is that opinion updates consist in averaging each opinion with that of its leader (in contrast to averaging each opinion over the set of opinions in a ball around it as in the Hegselmann-Krause dynamics). Also, the dynamics is extended to a particle system setting [45], namely to a stationary point process of $\mathbb{R}^{d}$ (in contrast to the Hegselmann-Krause dynamics which is only discussed on compact domains so far).
One of the surprising facts is that this simple dynamics, purely determined by the relative positions of agents, leads to a large collection of phenomena to which one can associate natural names in view of what is observed in social networks. Here are some instances of the phenomena observed in this dynamics in the Poisson case and that are discussed in detail in Section 3:
  • One agent can be a leader to several agents. The maximal number depends on the dimension. In dimension two, for all point configurations, a point can be a leader for up to 6 followers. This number is at most 5 a.s. in the Poisson case. This is related to the kissing number [22].
  • Each initial party has a.s. exactly one ”loop”, i.e., only one pair of mutually closest neighbors who will be referred to as the ultimate leader pair of the tree.
  • If two points form an ultimate leader pair, then, after one step, their opinions merge and will never change.
  • At each step of the dynamics, an agent either keeps the same leader or follows another agent. Similarly, a leader at any step can keep its followers, gain, or lose some followers.
  • Surprisingly, two agents in a follower/leader pair can swap their roles in one step of the dynamics. This follower inversion is reminiscent of the mentor/mentored scenario where the direction of leadership changes as time evolves.
  • As a result of this dynamics, at each step, agents either stay in their party or switch positions within a party, or change party. Equivalently, under a step of the dynamics, a party can either stay the same, split, loose, or gain new agents.
  • Some configurations of points lead to a party break, i.e., a situation where a point leaves one party and becomes a member of another one.
The identification of these purely geometry based dynamical phenomena is complemented by analytical and simulation results on the Poisson model. The paper starts with Section 2 where we give the notation and the point process background used in the rest of the document.
The first probabilistic question studied in the paper is about the spatial frequency of the phenomena described above at any given step. Section 4, we show that in the Poisson case, the frequencies in question admit representations in terms of integral geometry formulas involving integration domains which are semi-algebric sets of the Euclidean space and integrands which are determined by the volume of certain unions of balls. Section 5 is a complement to Section 4 which discusses numerical methods to evaluate the integral geometry formulas alluded to above and spatial simulation methods for the complementary estimation of the frequencies in question.
The second question we study is does the Poisson Follower graph of order $n$ have parties which are all of finite size a.s.? We prove this property for step 0 and step 1 of the dynamics. This result leverages the theory of dependent percolation over the square $\mathbb{Z}^{2}$ lattice and the mass transport principle. The necessary percolation theory background and the proof are presented in Section 6.
The third question studied in the paper is of dynamical system nature and concerns the limiting behavior of the Poisson follower point process of order $n$ and of the Poisson follower graph of order $n$, as $n$ tends to infinity. These questions are discussed in Section 7. We conjecture the classification of the points of the initial Poisson point process into two categories: ultimate followers, which remain followers forever and whose opinions only converge weakly to their limiting value, and utimate leaders whose opinion converges in total variation to their limiting value, which is hence reached in a finite number of steps. In the same section, we study stable trees, i.e., the trees that do not change their structure with time.
2 The Follower Model and its Background
2.1 Overview of Point Process
Let $||\bullet||$ denote the Euclidean norm on $\mathbb{R}^{2}$ and $B(A,r)$ the open ball of radius $r$ and center $A$.
Let $\Lambda$ be a $\mathbb{R}^{2}$ space equipped with its Borel $\sigma$- algebra $\mathcal{B}$ and Radon measure $\nu$. A point process $\Phi$ on $\Lambda$ is a random locally finite subset of $\Lambda$. One can also view $\Phi$ as a random counting measure on $\Lambda$, having the form $\Phi=\sum_{i\in\mathbb{N}}\delta_{S_{i}}$, where $\{S_{i}\}_{i\in\mathbb{N}}$ is a countable collection of points in $\Lambda$ with no accumulation points.
A point process is called simple if $\Phi(\{x\})\leqslant 1$ for all $x\in\Lambda$. A point process whose distribution is invariant under translations is called stationary. The intensity measure of a point process $\Phi$ is the measure on $\Lambda$ defined by
$\lambda(B)=\mathbb{E}[\Phi(B)],\,\,B\in\mathcal{B}(\Lambda).$
Definition 1.
For any counting measure $\mu=\sum_{i\in\mathbb{N}}\delta_{x_{i}}$ and $n\in\mathbb{N}$, its $n$-th power in the sense of products of measures is
$\mu^{n}=\sum_{(i_{1},...,i_{n})\in\mathbb{Z}^{n}}\delta_{(x_{i_{1}},...,x_{i_{n}})}.$
Define the $n$-th factorial moment measure as the following counting measure on $(\mathbb{R}^{2})^{n}$.
$\mu^{(n)}=\sum_{(i_{1}\neq...\neq i_{n})}\delta_{(x_{i_{1}},...,x_{i_{n}})}.$
Additionally, moment measures are an important object because they give information on the distribution of a point process.
Definition 2.
Moment measures. For a point process $\Phi$ on $\mathbb{R}^{2}$, let $\Phi^{n}$ be the $n$-th power of $\Phi$ and $\Phi^{(n)}$ be the $n$-th factorial moment of $\Phi$. We call $M_{\Phi^{n}}$ the $n$-th moment measure (the first moment measure is the mean measure) of $\Phi$ and $M_{\Phi^{(n)}}$ the $n$-th factorial moment measure.
Moment measures give important average structural characteristic of the process, such as level of clustering or repulsion.
Definition 3.
(Poisson point process) Let $\Lambda$ be a locally finite measure on l.c.s.h. space $\mathbb{G}$. A point process $\Phi$ is said to be Poisson with intensity measure $\Lambda$ if for all pairwise disjoint sets $B_{1},...,B_{j}\in\mathcal{B}(\mathbb{G})$, the random variables $\Phi(B_{1}),...,\Phi(B_{j})$ are independent Poisson random variables with respective means $\Lambda(B_{1}),...,\Lambda(B_{j})$; i.e. $\forall m_{1},...,m_{j}\in\mathbb{N}$,
$\mathbb{P}(\Phi(B_{1})=m_{1},...,\Phi(B_{j})=m_{j})=\prod_{i=1}^{j}\frac{\Lambda(B_{i})^{m_{i}}}{m_{i}!}e^{-\Lambda(B_{i})}.$ (1)
Definition 4.
(Homogeneous Poisson point process on $\mathbb{R}^{2}$) If $\Phi$ is a Poisson point process on $\mathbb{R}^{2}$ with intensity measure $\Lambda(dx)=\lambda\times dx$ where $\lambda\in\mathbb{R}^{*}_{+}$ and $dx$ denotes the Lebesgue measure, then $\Phi$ is called a homogeneous Poisson point process of intensity $\lambda$.
Throughout the paper, we denote by $\Phi=\sum_{i\in\mathbb{Z}}\delta_{x_{i}}$ the homogeneous Poisson point process on $\mathbb{R}^{2}$ with intensity $\lambda$, which serves as initial condition to the dynamics. For all $n\geqslant 0$, the Poisson Follower point process of order $n$ will be denoted by $\Phi_{n}$. Note that $\Phi=\Phi_{0}$.
For two points $x,y\in\Phi$, we denote by $B(x\rightarrow y)$ the open ball with center $x$ and radius $d(x,y)$, which is the distance between $x$ and $y$.
Palm Theory
Informally, the Palm measures of a point process $\Phi$ at a point $x\in\Lambda$ is the probability measure of $\Phi$ conditioned on having a point at location $x$. For a more detailed discussion on the matter, see [7, 38].
2.2 Notation for the Follower Dynamics
Consider a counting measure on $\mathbb{R}^{2}$. Take two points $A$ and $B$ in this counting measure. If the closest point to $A$ is $B$, we say that $A$ follows $B$. Other equivalent statements are, ”$A$ is a follower of BB”, or ”$B$ is a leader of AA”, and ”$B$ is followed by AA”, etc. We will also use notation $A\rightarrow B$ to mean that $A$ follows $B$. If $B$ follows $C$, and $A$ follows $B$, we call $C$ a leader of order two of $A$. The leader of $A$ will be denoted by $\mathscr{A}(A)=B$. The leader of order $k$ of $A$ will be denoted by $\mathscr{A}^{k}(A)$. Two agents A and B are part of the same follower chain if there is a $k$ such that $\mathscr{A}^{k}(A)=B$, or $l$ such that $\mathscr{A}^{l}(B)=A$. Note that this is an equivalence relation. We call ultimate leader pair agents that follow each other.
Because of scale invariance of the Poisson point process, we will assume without loss of generality that $\lambda=1$ in subsequent sections.
In the next Section we give the notation and the point process background used in the rest of the paper.
3 Phenomena Present in the Follower Dynamics
In this section we give some basic definitions to be used throughout the paper. We also define further the phenomena of interest introduced above and give some illustrations.
Forward and Backward Sets
We denote by $For(x,\Phi)$ the leader set of all orders of $x$ in $\Phi$, and we will call it the forward set of $x$. $Back(x,\Phi)$ denotes the follower set of all orders of $x$ in $\Phi$, and we call it the backward set of $x$. See Figure 2.
Figure 2: Representation of the forward and backward sets of an agent. The forward set of $x$ is connected to $x$ with dark gray arrows and the backward set of $x$ is with light gray ones. The agents connected by black arrows are in the follower party of $x$ but not in the backward or forward set of $x$.
Follower party
We call follower party any set of points belonging to one connected component of the follower graph. Namely, two points $x$ and $y$ belong to the same follower party if there exists a point $z$ (not necessarily different from $x$ or $y$), such that both $x$ and $y$ belong to the backward set of $z$, i.e., $x,y\in Back(z,\Phi)$.
One example can be found in Figure 2. Since the initial agent distribution is Poisson at step 0, every follower party is finite almost surely at step 0 [33].
Looking from any agent (point) of the system, here is a classification of situations that can happen to this agent when applying the dynamics.
”Out” edges
The ”out” edge of this agent or point, is either kept (leader keep) or changed (leader swap). Notice that these two situations are mutually exclusive since each point gets only one leader at any given time. For example in Figure 3, Agent 3 swapped its leader from 1 to 2, as a result of the dynamics.
Figure 3: Left: Example of a Follower Loss/Follower Gain. The initial positions are shown in black. The step 1 is in red, with the new positions denoted with the prime ’. Right: Example of formation of an ultimate leader pair of order 1. Same convention as in the left image
”In” edges
For the ”in” edges of the point, one can have a follower loss, a follower gain or a follower keep. Notice that in this case, all the situations can happen at the same time, since a point can have several followers. In the left Figure 3, Agent 1 lost its follower, Agent 3. Agent 2 gained a follower, namely Agent 3. Agent 1 kept Follower 2.
Ultimate leader pairs
A special case is when a new pair of ultimate leaders is formed. An example of this is shown in the right Figure 3. Let us call initial ultimate leader pairs, ultimate leader pair of order 0. New ultimate leader pairs, forming at step 1 will be called ultimate leader pairs of order 1, etc.
Figure 4: Example of a follower inversion.
Figure 5: Example of party fission
Follower inversion
Another special case is that when, after one step, the leader of an agent becomes its follower. We call such a situation a follower inversion. We do not count situations when the leader and follower become an ultimate leader pair at step 1 as a follower inversion. An example of follower inversion is shown in Figure 3. The initial positions are shown in black. The positions at the step 1 are shown in red, with the new positions denoted with the prime, like for example $1^{\prime}$. Agent 3 was following agent 2 initially, but at step 1, agent 2 follows agent 3, and agents 3 and 4 form an ultimate leader pair.
Looking at a party level, several situations can happen.
Party fission
The situation when there is a new ultimate leader is pair created within a party will be called a party fission. Figure 5 gives an example of a party fission. When a party increases its cardinality due to a swap of an agent from another party, we speak of party gain. The opposite event, i.e., a decrease in cardinality due to a swap of an agent from a party toward another preexisting party is called a party loss. A party can have both losses and gains at a given step.
Party restructuring
By party restructuring, we mean a situation where there is a swap in the party which does neither lead to a fission nor to a loss in the party, like the situation illustrated in Figure 3 right. A stable party is a party in which there are no swaps (and no losses or gains) at any future time step. Stable parties are discussed in Section 7.1.
Figure 6: Example of a 4 body swap. A’,B’, $A_{1}^{\prime}$ and $B_{1}^{\prime}$ are the positions of agents in the next time step. Positions in the second time step are denoted by AA" and $A_{1}$" and connections are shown in green
Party swap
The situation where an agent changes its follower party will be called a party swap. An example of a party swap is shown in Figure  6. Another example of a party swap is shown in Figure 7, where agent $z$ and its followers have a party swap.
Figure 7: Example of a party swap. Initial positions are shown in black and step 1 is in red. Agent $z$ swaps leaders and parties. Together with $z$, agents $x$ and $v$ also have a party swap.
4 body swap
In order to analyze the frequency of party swaps, it helps to focus on the necessary conditions for one agent to leave its party. There has to be an agent that is swapping to a leader from another party in order for the party swap to occur. See Figure 7 as an example of a party swap. In particular, for an agent $z$ to swap leaders, we need to be able to compare the distances between the agent $z$ and its leader $y$, and the distance between the agent $z$ and the potential leader at step 1, $u$. Thus we need to know the positions of the agent and the two swapping leaders at step 1, i.e., whether $d(z,u,\Phi_{1})<d(z,y,\Phi_{1})$? Therefore, we define a 4 body swap to be a leader swap that involves 4 different agents: $A,B,A_{1},$ and $B_{1}$, with, $A$ following $B$, and $A_{1}$ following $B_{1}$ at step 0. Then, at step 1, $A$ follows $A_{1}$ or $A_{1}$ follows $A$, i.e., a leader swap occurs. However, in order to know whether $A$ follows $A_{1}$ at step 1, we need to know where $B$ and $B_{1}$ moved to. For the analysis, we also need the positions of the leaders of $B$ and $B_{1}$, $C$ and $C_{1}$, respectively. See Figure 6 for an illustration. Using the notation of Figure 6, we want $d(A^{\prime},A_{1}^{\prime})<d(A^{\prime},B^{\prime})$.
Figure 8: Left: Initial positions and connections are shown in blue. Step 1 is shown in red. Agent $y$ starts following $x$ after 1 step, i.e., we see an example of a follower inversion. Agents $x$ and $u$ become ultimate leader pair of order 1. Right: Example of a party swap, where $v$ swaps parties and agents $u$ and $x$ form an ultimate leader pair of order 1.
Phenomena are not excluding of each other.
Under the Follower Dynamics, in the Poisson case, we observe all the phenomena described in this section and it can happen that after one step, several events happen at the same time on a given set of agents. For example, on an agent level, in the left Figure 8, we observe both a follower inversion and the creation of an ultimate leader pair of order 1. On the party level, the left of Figure 8 also gives an example of simultaneous party restructuring and party fission since there are 2 pairs of ultimate leaders. On the agent level, in the right Figure 8 we observe both a party swap and the formation of an ultimate leader pair of order 1. The right Figure 8 is also an example of a simultaneous party restructuring and party swap.
4 Frequency of the Phenomena
In this Section, we develop a systematic method to calculate or estimate bounds on the spatial frequencies of related events using integral geometry. We demonstrate that this involves evaluating integrals of specific exponential functions over semi-algebraic sets. Semi-algebraic sets are subsets of $\mathbb{R}^{2}$, which are finite unions of solutions to polynomial equations and inequalities with coefficients in $\mathbb{R}$ [42]. This approach applies to each phenomenon listed above.
First, as a warm-up, in Subsection 4.1, we examine all pairs of points and show how to determine how often they are each other’s closest neighbors, i.e., we calculate the spatial density of ultimate leader pairs of order zero. This is a known result [46], but we derive it here in a way that can be extended to other phenomena. We also show how to calculate the density of ultimate leader pair of order one.
4.1 Integral Geometry Estimates of Densities of Ultimate Leader Pairs
In this subsection, we give an integral geometry representation of ultimate leader pairs and then give numerical estimates. We illustrate the method of calculating the densities by showing a step by step calculation of the density of ultimate leader pairs in the initial configuration, i.e., the density of order 0. One can calculate frequencies of other configurations in the same manner. Then, we calculate an upper bound and an exact integral geometry formula for the density of ultimate leaders of order 1, type 1 or density of order one type 1. Again, by ultimate leader pairs we mean a nearest neighbor pair.
4.1.1 Order Zero
Consider the point process $N^{(2)}$ consisting of pairs of points of $\Phi$ which are mutually closest points. Namely
$N^{(2)}=\sum_{i\neq j\in\mathbb{Z}}\delta_{x_{i},x_{j}}1_{\Phi(B(x_{i}\rightarrow x_{j}))=1},\\ 1_{\Phi(B(x_{j}\rightarrow x_{i}))=1},$
where $\Phi(B(x_{j}\rightarrow x_{i}))$ denotes the number of points in the open ball centered at $x_{j}$ with the radius $d(x_{j},x_{i})$.
The mean measure of $N^{(2)}$ is given by, for $A\subset\mathbb{R}^{2}\times\mathbb{R}^{2}$,
$\begin{split}\mathbb{E}[N^{(2)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y\in A}1_{x\neq y}1_{\Phi(B(x\rightarrow y))=1}1_{\Phi(B(y\rightarrow x))=1}\Phi^{(2)}(dx\times dy)],\end{split}$ (2)
where $\Phi^{(2)}$ is the Poisson factorial moment measure of order 2 ([7] Section 3.3.2). By the higher order Campbell-Little-Mecke formula [5]
$\begin{split}\mathbb{E}[N^{(2)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y\in A}\mathbb{E}^{x,y}[1_{\Phi(B(x\rightarrow y))=1}1_{\Phi(B(y\rightarrow x))=1}]\lambda^{(2)}(dxdy),\end{split}$ (3)
where $\mathbb{E}^{x,y}$ is the two point Palm expectation and $\lambda^{(2)}$ is the factorial Poisson moment measure of order 2 ([7] Section 3.3.2). Recall that for a Poisson point process, the $n^{th}$ factorial moment measure equals the $n^{th}$ power of the intensity measure. We will be using this over and over again in the calculations that follow.
Now by Slivnyak’s theorem [5]
$\mathbb{E}[N^{(2)}(A)]=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}1_{x,y\in A}\mathbb{E}[1_{\Phi^{\prime}(B(x\rightarrow y))=1}1_{\Phi^{\prime}(B(y\rightarrow x))=1}]\lambda^{2}dxdy,$ (4)
where $\Phi^{\prime}=\Phi+\delta_{x}+\delta_{y}$. Using the change of variables $(x,y)\rightarrow(x,u)$ with u=yxu=y-x and taking $A=C\times\mathbb{R}^{2}$, with $C\in\mathcal{B}(\mathbb{R}^{2})$, we get
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}1_{x\in C}1_{u\in\mathbb{R}^{2}}\mathbb{E}[1_{\Phi^{\prime}\circ\theta_{x}(B(o\rightarrow u))=1}1_{\Phi^{\prime}\circ\theta_{x}(B(u\rightarrow 0))=1}]\\ \lambda^{2}dxdu.$ (5)
The expectation is the probability for point $u$ to be the closest to 0 and for 0 to be closest to $u$. Then
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{\mathbb{R}^{2}}1_{u\in\mathbb{R}^{2}}\mathbb{E}[1_{\Phi(B(o\rightarrow u))=0}1_{\Phi(B(u\rightarrow 0))=0}]\\ \lambda^{2}du.$ (6)
Because of symmetry, and after switching to polar coordinates, we get:
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{\theta\in[0,2\pi]}\int_{\mathbb{R}}ve^{-\lambda v^{2}\pi}e^{-\lambda v^{2}(\pi/3+\sqrt{3}/2)}\lambda^{2}dv.$ (7)
Hence
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{\mathbb{R}}2\pi ve^{-\lambda v^{2}\pi}e^{-v^{2}(\pi/3+\sqrt{3}/2)}\lambda^{2}dv.$ (8)
To evaluate the integral, we use the change of variables $v\rightarrow r$ with $r=\pi v^{2}$, and recall that $\lambda=1$. Then we get
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})])=|C|\int_{0}^{\infty}e^{-r(\frac{\pi+(\pi/3+\sqrt{3}/2)}{\pi})}dr$
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})])=|C|\cdot\frac{\pi}{\pi+\pi/3+\sqrt{3}/2}\approx 0.62|C|.$
So the density of the ultimate leader pair of order 0 is $0.62$.
4.1.2 Density of Ultimate Leader Pairs of Order One
In this subsection, we continue with similar calculations to determine the intensity of ultimate leader pairs of order $1$. First, we provide an upper bound on this frequency and then explain how to adjust it to find the exact frequency. Each calculation requires the positions of at least $4$ points to know the positions at step 1. There are two distinct configurations in which we obtain ultimate leader pairs at step 1. The first involves situations where two followers of one point become ultimate leader pairs at step 1. The second involves cases where a leader and a follower become an ultimate leader pair at the next step. Some examples of each are shown in Figure 9. Frequencies of both types are computed similarly, so we focus on the detailed calculation of type 1 only.
Figure 9: Left: Four points and their positions at step 1. Example of ultimate leader pair of order 1, type 1: two followers of one point become an ultimate leader pair at step 1. Right: Example of an ultimate leader pair of order 1, type 2. A point and its leader become an ultimate leader pair at step 1.
Type 1
As mentioned above, ultimate leader pairs of order 1, type 1 are formed from the configurations where the two step 0 followers of one agent become an ultimate leader pair at step 1. We first give an upper bound on the frequency of type 1 and then we analyze the exact frequency.
Let $\mathcal{D}\in{\mathcal{B}}^{4}$ be the set of all distinct points $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ that satisfy the conditions (9)-(12) defined below:
$d(x_{1},x_{3})<d(x_{1},x_{2}),\,d(x_{1},x_{3})<d(x_{1},x_{4}),$ (9)
are the conditions needed for $x_{1}$ to follow $x_{3}$, in the absence of other points than these four points;
$d(x_{2},x_{3})<d(x_{2},x_{1}),\,d(x_{2},x_{3})<d(x_{2},x_{4}),$ (10)
are the conditions needed for $x_{2}$ to follow $x_{3}$ as well, also in the absence of other points. Finally
$d(x_{3},x_{4})<d(x_{3},x_{1}),\,d(x_{3},x_{4})<d(x_{3},x_{2}),$ (11)
are the conditions needed for $x_{3}$ to follow $x_{4}$ under the same conditions. The phenomenon we are interested in is that where, in the above configuration,
$d(x_{1}^{\prime},x_{2}^{\prime})<d(x_{1}^{\prime},x_{3}^{\prime}),\,d(x_{2}^{\prime},x_{1}^{\prime})<d(x_{2}^{\prime},x_{3}^{\prime}),$ (12)
where $x_{1}^{\prime}=\frac{x_{1}+x_{3}}{2}$, $x_{2}^{\prime}=\frac{x_{2}+x_{3}}{2}$, and $x_{3}^{\prime}=\frac{x_{3}+x_{4}}{2}$. That is $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are mutual closest neighbors in the absence of other points. Note that $\mathcal{D}$ is a semi-algebraic set [42].
For the points $(x_{1},...,x_{4})$ to be involved in the formation of an ultimate leader pair of type 1, it is necessary but not sufficient that $(x_{1},\ldots,x_{4})$ belong to $\cal D$}. For this to happen, in addition, it must be that there are no other points of the Poisson P.P. that change the facts that both $x_{1}$ and $x_{2}$ follow $x_{3}$, $x_{3}$ follows $x_{4}$, and $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are mutually nearest neighbors.
Let $\Phi_{1,2,3,4}$ denote the point process $\Phi$ restricted to the set $x_{1},x_{2},x_{3}$, and $x_{4}$. Let $N_{1}^{(4)}$ be the point process of quadruples of points of the Poisson P.P. $\Phi$ that belong to $\cal D$}, and are such that the following event $M_{1}$ holds: in $\Phi$, both $x_{1}$ and $x_{2}$ follow $x_{3}$ and $x_{3}$ follows $x_{4}$.
In a first step, we evaluate the spatial frequency $\beta^{1}$ of the event $M_{1}$, which is an upper bound on the frequency of ultimate leader pairs of order 1, type 1.
For $A\in\mathcal{B}^{4}$, the mean measure of $N_{1}^{(4)}$ is given by,
$\mathbb{E}[N_{1}^{(4)}(A)]=\mathbb{E}[\sum_{x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}],$
where $\Phi$ is the P.P.P. and $B(x\rightarrow y)$ is the open ball of center $x$ and radius |xy||x-y|. In integral form, this is
$\mathbb{E}[N_{1}^{(4)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}$
$1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}\Phi^{(4)}(dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].$
By the higher order Campbell-Little-Mecke formula,
$\mathbb{E}[N_{1}^{(4)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}$
$\mathbb{E}^{x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}]\lambda^{(4)}(dx_{1}dx_{2}dx_{3}dx_{4}),$
where $\mathbb{E}^{x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x_{1},...,x_{4}$, and $\lambda^{(4)}$ is the Poisson factorial moment measure of order 4. By Slivnyak’s theorem,
$\begin{split}\mathbb{E}[N^{(4)}_{1}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}]\lambda^{4}dx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
where $\hat{\Phi}=\Phi+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P of intensity $\lambda$. We can simplify the expression a bit more using stationarity of the Poisson point process. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C\in\mathcal{B}(\mathbb{R}^{2})$ a compact. Because we have a stationary point process,using the change of variables, $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
$\mathbb{E}[N_{1}^{(4)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{\cal D}}e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0))}\lambda^{4}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4}=|C|\beta_{1},$
with $\tilde{{\cal D}}=\{(x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}\}$.
This is the announced integral over a semi-algebraic set. Note that this integral is twice the frequency of interest since we do not distinguish whether $d(x_{1},x_{3})<d(x_{2},x_{3})$ or the other way around. In order to evaluate the integral, we need to write a formula for the volume of the union of 3 balls. This is discussed in Section 5.2.
Now that we have all the elements, we have a formula to calculate exact value of the volume of the union of balls $\mathbb{U}$. For compactness, the integral of the frequency of type 1 can be written as:
$\mathbb{E}[N_{1}^{(4)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\lambda^{4}1_{\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in{\cal D}_{0}}e^{-\lambda\mathbb{U}}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
with ${\cal D}_{0}={\cal D}\cap\{x_{3}=0\}$.
We now discuss the exact calculation. In order to get the event of interest, one should in addition have $\Phi_{1}(B(x_{1}\to x_{2}))=1$ and $\Phi_{1}(B(x_{2}\to x_{1}))=1$, with $\Phi_{1}$ the point process at step 1. In other words, certain refinements of the above configurations should be removed from the counting. We can order these refinements in the disjoint and exhaustive categories listed below
  1. 1.
    There is an extra point $x$ in $\Phi$ that follows $x_{1}$ in $\Phi$ and such that the distance from $x^{\prime}$ to $x_{1}^{\prime}$ is less than that from $x_{1}^{\prime}$ to $x^{\prime}_{2}$. The case where there is an extra point $x$ in $\Phi$ that follows $x_{2}$ in $\Phi$ and such that the distance $d(x,x_{2},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$ is analogous and is counted here as well. This is due to a fact that we do not distinguish whether $d(x_{1},x_{3})<d(x_{2},x_{3})$ or vice versa.
    For evaluating the frequency of this event, we have to enrich $\cal D$} by adding $x$ and state that $x$ follows $x_{1}$ and that none of the points $x_{1},x_{2},x_{3},x_{4}$ follows $x$, that is
    $\begin{split}d(x,x_{1})<d(x,x_{2}),\,d(x,x_{1})<d(x,x_{3}),\,d(x,x_{1})<d(x,x_{4}),\\ d(x_{1},x_{3})<d(x_{1},x),\,d(x_{2},x_{3})<d(x_{2},x),\,d(x_{4},x_{3})<d(x_{4},x).\end{split}$ (13)
    This adds 6 quadratic inequalities. Finally, the condition that $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$, gives one more quadratic inequality
    $d(x^{\prime},x_{1}^{\prime})<d(x^{\prime}_{1},x^{\prime}_{2}),$ (14)
    where $x^{\prime}=\frac{x+x_{1}}{2}$. So the frequency of this refinement can also be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{1}$, which is a refinement of $\cal D$} with one more variable and 7 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{5}$, the mean measure of $N_{1,1}^{(5)}$, which is the point process of the 5-tuples of points satisfying the above conditions, is given by,
    $\mathbb{E}[N_{1,1}^{(5)}(A)]$ $=$ $\mathbb{E}[\sum_{x,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}],$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,1}^{(5)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}\\ 1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}\Phi^{(5)}(dx\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    By the higher order Campbell-Little-Mecke formula,
    $\mathbb{E}[N_{1,1}^{(5)}(A)]$ $=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}$
    $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}]\lambda^{(5)}(dxdx_{1}dx_{2}dx_{3}dx_{4}),$
    where $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x,x_{1},...x_{4}$, and $\lambda^{(5)}$ is the factorial Poisson moment measure of order 5. By Slivnyak’s theorem,
    $\begin{split}\mathbb{E}[N_{1,1}^{(5)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}\\ 1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}1_{\hat{\Phi}(B(x\rightarrow x_{1}))=1}]\lambda^{5}dxdx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
    where $\hat{\Phi}=\Phi+\delta_{x}+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P with intensity $\lambda$. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C$ a compact. Because we have a stationary point process, using the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,1}^{(5)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{1}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow\tilde{x}_{1})))}\lambda^{5}d\tilde{x}\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with $\tilde{{\cal D}_{1}}=\{(x-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x,\,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{1}\}$.
    Similar to the derivation for the union of 3 balls, one can calculate the volume of the union of 4 balls.
  2. 2.
    There is an extra point $x$ in $\Phi$ that follows $x_{3}$ in $\Phi$ and such that the distance $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$. The case where $d(x,x_{2},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$ is symmetric.
    For evaluating the frequency of this event, we have to enrich $\cal D$ by adding $x$ and state that $x$ follows $x_{3}$ and that none of the points $x_{1},x_{2},x_{3}$ follow $x$.
    $\begin{split}d(x,x_{3})<d(x,x_{1}),\,d(x,x_{3})<d(x,x_{2}),\,d(x,x_{3})<d(x,x_{4}),\\ d(x_{1},x_{3})<d(x_{1},x),\,d(x_{2},x_{3})<d(x_{2},x),\,d(x_{4},x_{3})<d(x_{4},x).\end{split}$ (15)
    This adds 5 quadratic inequalities. Finally, the condition that $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$, gives one more quadratic inequality:
    $d(x^{\prime},x_{1}^{\prime})<d(x^{\prime}_{1},x^{\prime}_{2}),$ (16)
    where $x^{\prime}=\frac{x+x_{1}}{2}$. So the frequency of this refinement can be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{2}$, which is a refinement of $\cal D$} with one more variable and 6 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{5}$, the mean measure of $N_{1,2}^{(5)}$, which is the point process of the 5-tuples of points satisfying the conditions above, is given by,
    $\mathbb{E}[N_{1,2}^{(5)}(A)]$ $=$ $\mathbb{E}[\sum_{x,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}],$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,2}^{(5)}(A)]&=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1},\\ &1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}\Phi^{(5)}(dx\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    By the higher order Campbell-Little-Mecke formula,
    $\begin{split}\mathbb{E}[N_{1,2}^{(5)}(A)]&=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}\\ &1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}]\lambda^{(5)}(dxdx_{1}dx_{2}dx_{3}dx_{4}),\end{split}$
    where $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x,x_{1},...x_{4}$, and $\lambda^{(5)}$ is the factorial Poisson moment measure of order 5. By Slivnyak’s theorem,
    $\begin{split}\mathbb{E}[N_{1,2}^{(5)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{0}^{\prime}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}\\ 1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}1_{\hat{\Phi}(B(x\rightarrow x_{3}))=1}]\lambda^{5}dxdx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
    where $\hat{\Phi}=\Phi+\delta_{x}+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P with intensity $\lambda$. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C$ a compact. Because we have a stationary point process, using the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,2}^{(5)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{2}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow 0)))}\lambda^{5}d\tilde{x}\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with $\tilde{{\cal D}_{2}}=\{(x-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x,\,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{2}\}$.
    Again, one can calculate the volume of the union of 4 balls.
  3. 3.
    There are two extra points $x$ and $y$ such that $x$ follows $y$ (4 inequalities) and none of the points $x_{1},\ldots,x_{3}$ follow either $x$ or $y$ (6 inequalities). Thus we have
    $d(x,y)<d(x,x_{1}),\,d(x,y)<d(x,x_{2}),\,d(x,y)<d(x,x_{3}),\,d(x,y)<d(x,x_{4}),$
    $d(x_{1},x_{3})<d(x_{1},x),\,d(x_{1},x_{3})<d(x_{1},y),$
    $d(x_{2},x_{3})<d(x_{2},x),\,d(x_{2},x_{3})<d(x_{2},y),$
    $d(x_{3},x_{4})<d(x_{3},x),\,d(x_{3},x_{4})<d(x_{3},y).$
    WLOG take $x^{\prime}=\frac{x+y}{2}$ to be closer to $x^{\prime}_{1}$ than $x^{\prime}_{1}$ to $x^{\prime}_{2}$, i.e.,
    $d(x^{\prime}_{1},x)<d(x^{\prime}_{1},x^{\prime}_{2}).$
    This amounts to 2 more variables and 11 more quadratic inequalities and one more empty ball conditions ($\Phi(B(x\to y))=1$). So the frequency of this refinement can be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{3}$, which is a refinement of $\cal D$} with two more variables and 11 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{6}$, the mean measure of $N_{1,3}^{(6)}$, which is the point process of the 6-tuples of points satisfying the conditions above, is given by
    $\mathbb{E}[N_{1,3}^{(6)}(A)]$ $=$ $\mathbb{E}[\sum_{x,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{3}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow y))=1}],$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,3}^{(6)}(A)]&=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{3}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1},1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}\\ &1_{\Phi(B(x\rightarrow y))=1}\Phi^{(5)}(dx\times dy\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    By the higher order Campbell-Little-Mecke formula,
    $\begin{split}\mathbb{E}[N_{1,3}^{(6)}(A)]&=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{3}}\mathbb{E}^{x,y,x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}\\ &1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow y))=1}]\lambda^{(6)}(dx\,dy\,dx_{1}dx_{2}dx_{3}dx_{4}),\end{split}$
    where $\mathbb{E}^{x,y,x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x,y,x_{1},...,x_{4}$, and $\lambda^{(6)}$ is the factorial Poisson moment measure of order 6. By Slivnyak’s theorem,
    $\begin{split}\mathbb{E}[N_{1,3}^{(6)}(A\cap{\cal D}_{3})]&=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{0}^{\prime}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}\\ &1_{\hat{\Phi}(B(x\rightarrow y))=1}]\lambda^{5}dx\,dy\,dx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
    where $\hat{\Phi}=\Phi+\delta_{x}+\delta_{y}+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P with intensity $\lambda$. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C$ a compact. Because we have a stationary point process, using the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{y}=y-x_{3}$ $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,3}^{(6)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]$ $=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{3}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow\tilde{y})))}\lambda^{6}d\tilde{x}d\tilde{y}\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with
    $\tilde{{\cal D}_{3}}=\{(x-x_{3},\,y-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},0,x_{4}-x_{3}):(x,y,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{3}\}.$
Let $\beta_{i}$ the frequency of event $i=1,2,3$ in the above list. Then the frequency of interest is
$\frac{1}{2}(\beta^{1}-\sum_{i=1}^{3}\beta_{i}).$ (17)
We use numerical methods to evaluate all the frequencies. Numerical estimation is discussed in the next Section.
5 Numerical Estimation Methods
We use two different methods to estimate our frequency calculations. The first is discrete event spatial simulation, which depends on sampling point processes and exploiting the ergodic properties of the Poisson point process and related factors of such point processes [66, 7]. This approach is grounded in the ergodic theory of point processes theory [66]. The second method involves numerical evaluation of integrals based on semi-algebraic sets discussed above. We will explain both methods in detail and provide estimated values for the frequencies mentioned earlier. For each, we will first provide some theoretical background and then describe how we apply it in our evaluation. For simulation, we include 95% confidence intervals for all frequencies discussed in previous sections. Finally, we compare the results obtained from both.
5.1 Discrete Event Spatial Simulation
The theoretical background about ergodicity discussed in this section comes from [7].
Definition 5.
[7] A stationary framework $(\Omega,\mathcal{A},\{\theta_{t}\}_{t\in\mathbb{R}^{d}},\mathbb{P})$ is said to be ergodic if
$\lim_{a\rightarrow\infty}\frac{1}{(2a)^{d}}\int_{[-a,a]^{d}}\mathbb{P}(A_{1}\cap\theta_{-x}A_{2})dx=\mathbb{P}(A_{1})\mathbb{P}(A_{2}),\,\,\forall A_{1},A_{2},\in\mathcal{A}.$
It is said to be mixing if
$\lim_{|x|\rightarrow\infty}\mathbb{P}(A_{1}\cap\theta_{-x}A_{2})=\mathbb{P}(A_{1})\mathbb{P}(A_{2}),\,\,\forall A_{1},A_{2}\in\mathcal{A}.$
Lemma 1.
For all $n$, the follower point process of order $n$ is a factor of a Poisson point process, and hence is mixing.
Discrete event spatial simulation starts with sampling a Poisson point process in the fixed bounded window $W\subset\mathbb{R}^{2}$ (the observation window). We pick a convex $W$, typically a square. In order to simulate a homogeneous Poisson point process, we use the ideas from [66]. In [66] it is suggested that simulating in a compact region $W$ can be split into two parts. Namely, the number of points in $W$ is determined from the simulation of a Poisson random variable. Then the positions of points in $W$ are obtained from simulating a binomial point process with that number of points. Since we are interested in the point process in the whole of $\mathbb{R}^{2}$, we take $W$ big, hence there are small boundary effects.
Then we apply the Follower Dynamics to all the points simultaneously. We still have mixing properties satisfied. All simulations are done using MATLAB. For each agent, we keep track of its leader at every step. In order to calculate any frequency, we count the number of agents in the window for which our sets of conditions are satisfied and we divide by the total number of agents in the window.
For example for the frequency of ultimate leader pairs of order 0, we check for all pairs of agents and count the number of mutually closest neighbors at step 0. Since we keep track of the leader for each agent, we just check whether two agents are each others’ leaders.
Similarly, in order to determine the frequency of the ultimate leader pairs of order 1 type 1, we check for all pairs of agents, whether they had the same leader (but are not each others’ leaders) at step 0 and are each others’ leaders at step 1.
For the frequency of the ultimate leader pairs of order 1 type 2, we check, for all pairs of agents, whether one was a leader of the other at step 0 and both are each others’ leaders at step 1.
Table 1 contains frequencies and $95\%$ confidence interval obtained using discrete event spatial simulation. In order to obtain a $95\%$ confidence interval we run simulation many times and keep track of the statistics. The statistics is obtained from $40$ samples containing in mean each 20000 points. We observe that the frequency of ultimate leader pairs of order 1 is in the interval $[0.021,0.023]$ with the $95\%$ confidence. Similarly, frequency of inversion is in the interval $[0.13,0.14]$ with the same confidence.
Frequency of $95\%$ Confidence interval
Ultimate leader pair of order 0 $[0.6203,0.6227]$
Order 1 type 1 $[0.011,0.012]$
Order 1 type 2 $[0.0100,0.0105]$
4 body swap $[0.000063,0.0001]$
Inversion Type 1 $[0.138,0.140]$
Inversion Type 2 $[0.00019,0.00027]$
Table 1: Frequencies and $95\%$ confidence intervals obtained using discrete event spatial simulation
5.2 Integral Equations and Semi-Algebraic Sets
When exploiting our integral geometry representations of frequencies, we end up with expressions that are integrals over unions of balls. We hence need to compute the volumes of the union of balls. First, we give some more background. In this subsection we first represent the unions of balls as a semi-algebraic set.
The setting is $\mathbb{R}^{d}$. Let $k$ be a positive integer. Let $c_{1},\ldots,c_{k}$ be arbitrary points of $\mathbb{R}^{d}$ which are the centers of balls of positive radii $r_{1},\ldots,r_{k}$, respectively. We are interested in the volume of
$\mathbb{U}:=\cup_{i=1}^{k}B(c_{i},r_{i}),$
with $B(c,r)$ the closed ball of center $c$ and radius $r$.
Let $\mathcal{S}$ be set of all non-empty subsets of $1,\ldots,k]$ For all $s\in S$, let
$\mathbb{V}_{s}=\cap_{i\in s}^{k}B(c_{i},r_{i})\cap_{j\notin s}^{k}B(c_{j},r_{j})^{c}.$
We have
$\mathbb{U}=\cup_{s\in S}\mathbb{V}_{s},$
where the last union is disjoint. Hence
$\mid\mathbb{U}\mid=\sum_{s\in S}\mid\mathbb{V}_{s}\mid.$
But for all $s$, $\mathbb{V}_{s}$ is the following semi-algebraic set:
[…]
Hence $\mathbb{U}$ is the disjoint union of the $2^{k}-1$ semi-algebraic sets $\{\mathbb{V}_{s},s\in S\}$ and its volume is the sum of the volumes of these semi-algebraic sets of $\mathbb{R}^{d}$.
Once we have a semi-algebraic set we need to compute an integral over it.
Numerical Result:
In practice the process of calculating these integrals is deterministic. However, the algorithm is quite slow $O(N^{6})$, where $N$ is the number of spacing. The algorithm for calculating the upper bound on the frequency of ultimate leader pairs of order 1, type 1 was written in MATLAB and a value of $0.03$ was obtained. Even this upper bound estimate takes several hours to run in order to obtain the approximate value of the integral.
6 The Follower Model and Percolation
This section presents our main results and proofs concerning the size of parties. We begin by providing a new proof that the Poisson Descending Chain is almost surely finite. Then, we extend this result to the Poisson Follower Model graph at step 1.
Let’s start with an overview of the ideas behind this proof. These ideas will apply to each step of the Follower Dynamics, with some technical differences at each stage.
We tessellate the plane with squares of side length $2a$, ensuring every point of the Follower Model lies within some square. Based on certain local conditions, we classify squares as open or closed. We show that the open squares percolate while the closed squares do not. Therefore, if an infinite party contains the origin, the forward set from 0 must cross an open square at some point. This proof mainly relies on the fact that distances decrease, which is why we analyze the forward set from 0. Once it encounters an open square, the forward set cannot ”travel” too far. Specifically, it can cross only finitely many open and closed squares. As a result, the forward set must be finite a.s. Then, using the Point Map Classification [8], we conclude that the party is finite a.s..
6.1 Follower Parties at Step 0
We introduce here our novel proof that the nearest neighbor graph on a Poisson point process does not percolate.
Theorem 1.
The nearest neighbor graph on a Poisson point process does not percolate.
Proof.
The proof consists in first showing that the forward set of a typical point is integrable a.s. Then, using the Mass Transport Principle [7], we prove that the backward (or the follower set) is also finite a.s.. The proof relies on the property of the Follower Model that every agent always follows its nearest neighbor, so the distances are always decreasing along the forward path.
Let $\epsilon>0$ be fixed.
For a point $x\in\Phi$ define its contraction ratio $\rho(x)$ as
$\rho(x)=\frac{d(h(x,\Phi),h^{2}(x,\Phi))}{d(x,h(x,\Phi))}1_{h^{2}(x,\Phi)\neq x}.$ (18)
We exclude the nearest neighbor pairs from consideration for the ratio. For mutually closest neighbors, we define the contraction ratio to be 0.
Theorem 2.
Under $\mathbb{P}^{0}$ the forward and the backward sets of 0 are a.s. finite.
Proof.
The proof is based on a series of lemmas. Before we begin the proof: In this theorem, a path that ends with a mutually closest neighbor pair is considered infinite.
The forward path can end at finite distance with a mutually closest neighbor pair and the statement would be true, or it doesn’t end that way. If it does not, let us consider the following site percolation problem:
Define the square boxes to be of length $2a$ and center $2na$, $n\in\mathbb{Z}^{2}$. Declare a square open if it satisfies the following conditions:
  1. i
    It contains at least one point;
  2. $i$
    Each point in the square has its nearest neighbor at a distance at most $a$, where $2a$ is the side length of the squares;
  3. iii
    Each point in the square has a contraction ratio of at most $\rho$.
A square is closed if it is not open. In Lemma 2 we give an explicit calculation of the probability that a point has a contraction ratio greater than $1-\epsilon$, where $\epsilon$ is a constant we can vary. Then in Lemma 3, we show that the probability that a square is closed can be made arbitrarily close to 0 by choosing $a$ large enough and $\epsilon$ small enough.
The structure of the proof of Theorem 2 is divided into three separate components:
  1. 1.
    For a proper choice of lattice side length $2a$, there is a percolation of the open squares;
  2. 2.
    Each infinite path from 0 crosses infinitely many open squares;
  3. 3.
    Use mass transport formula to show that the forward path from 0 has to be finite and then show that the backward set from 0 also has to be finite a.s.
Lemma 2.
For all $a$, the Palm probability that the distance to the nearest point is less than $a$ and that the contraction ratio of the origin is greater than $1-\epsilon$ tends to 0 as $\epsilon\rightarrow 0$.
Proof.
We use the notation represented in Figure 10. Denote the distance from $x$ to its nearest neighbor $y$, by $r_{1}$. Let the distance $d(h(x,\Phi),h^{2}(x,\Phi))$ be $r_{2}$. For all $r_{2}<r_{1}$, let $A_{1}$ be the region $B(y,r_{2})\setminus B(x,r_{1})$, and $A_{2}$ be the region $B(y,r_{1})\setminus B(y,r_{2})$. For the contraction ratio to be more than $1-\epsilon$, we need region $A_{1}$ to be empty of other points, and $A_{2}$ to contain at least one point for $r_{2}=r_{1}(1-\epsilon)$.
Denote by $p_{0}$ the probability of the event that point $x$ has a nearest neighbor at distance $r_{1}$ and the neighbor of $x$ has its nearest neighbor at a distance between $r_{2}$ and $r_{1}$.
We are only interested in the case when $r_{1}<a$, so that the probability $p_{0}(a)$ can be written as follows:
$p_{0}(a,\epsilon)=\int_{0}^{a}\pi[\mbox{nearest neighbor is at distance }r_{1}<a]$
[…]
where $\pi$ is the Poisson nearest neighbor density, i.e., the Rayleigh density.
Figure 10: Contraction ratio $\rho(x)=r_{2}/r_{1}$. Gray shaded area depicts $A_{1}$ and red shaded area depicts $A_{2}$
In order to get explicit probabilities, we first need to calculate different areas. Let us denote by $A_{intersection}$, the area of the intersection of two balls, one centered at $x$ of radius $r_{1}$, and the other at $y$ of radius $r_{2}$.
$A_{intersection}=r_{2}^{2}\cos^{-1}\Big(\frac{r_{2}}{2r_{1}}\Big)+r_{1}^{2}\cos^{-1}\Big(\frac{2r_{1}^{2}-r_{2}^{2}}{2r_{1}^{2}}\Big)-\frac{1}{2}\sqrt{r_{2}^{2}((2r_{1})^{2}-r_{2}^{2})}.$
If $r_{1}=r_{2}$ then the area of intersection simplifies to $A_{eq}=r_{1}^{2}\big(\frac{2\pi}{3}-\frac{\sqrt{3}}{2})$.
The area of region $A_{1}$ is
$A_{1}=r_{2}^{2}\pi-A_{intersection}.$
Similarly, the area of region $A_{2}$ is
$A_{2}=(r_{1}^{2}\pi-A_{eq})-A_{1}=(r_{1}^{2}\pi-A_{eq})-r_{2}^{2}\pi+A_{intersection}.$
The total probability of a point having contraction ratio greater than $1-\epsilon$ is hence
$p_{0}(a,\epsilon)=\int_{0}^{a}2\lambda\pi r_{1}e^{-\lambda r_{1}^{2}\pi}\cdot e^{-\lambda A_{1}}\cdot(1-e^{-\lambda A_{2}})dr_{1},$
with $r_{2}=r_{1}(1-\epsilon)$. Since $\epsilon\ll 1$, throughout the calculations below we apply a first order Taylor expansions; with $(1-\epsilon)^{2}\approx 1-2\epsilon$ and $\sqrt{1+\epsilon}\approx 1+\frac{1}{2}\epsilon$.
Combining the equations for $A_{1}$, $A_{2}$ and $A_{eq}$ all together into $p_{0}$ we write it as a difference of two integrals
$p_{0}(a,\epsilon)=\int_{0}^{a}2\lambda\pi r_{1}e^{-\lambda r_{1}^{2}\pi}\cdot e^{-\lambda A_{1}}dr_{1}-\int_{0}^{a}2\lambda\pi r_{1}e^{-\lambda r_{1}^{2}\pi}\cdot e^{-\lambda(A_{1}+A_{2})}dr_{1}.$
Let $I_{1}=\int_{0}^{a}2\lambda\pi r_{1}e^{-\lambda r_{1}^{2}\pi}\cdot e^{-\lambda A_{1}}dr_{1}$ and $I_{2}=\int_{0}^{a}2\lambda\pi r_{1}e^{-\lambda r_{1}^{2}\pi}\cdot e^{-\lambda(A_{1}+A_{2})}dr_{1}$.
In calculations that follow, we again use $\lambda=1$.
We calculate each integral separately and then combine them together to find
$p_{0}(a,\epsilon)=I_{1}-I_{2}.$
First we can simplify the expression for $A_{1}$:
$A_{1}=r_{1}^{2}\Big((1-\epsilon)^{2}(\cos^{-1}(\frac{1-\epsilon}{2})-\pi)+\cos^{-1}(1-\frac{(1-\epsilon)^{2}}{2})-\frac{1-\epsilon}{2}\sqrt{4-(1-\epsilon)^{2}}\Big).$
Using a Taylor series expansion of first order, we can approximate $A_{1}$ as
$A_{1}\approx r_{1}^{2}\Big((1-2\epsilon+o(\epsilon))(-\frac{2\pi}{3}+\frac{\epsilon}{\sqrt{3}}+o(\epsilon))+\frac{\pi}{3}-\frac{9\epsilon}{8}-\frac{\sqrt{3}}{2}(1-\frac{2\epsilon}{3}+o(\epsilon))\Big),$
which simplifies to
$A_{1}\approx r_{1}^{2}\Big(-\frac{\pi}{3}-\frac{\sqrt{3}}{2}+\epsilon(\frac{4\pi}{3}-\frac{9}{8})+o(\epsilon)\Big).$
Thus $I_{1}$ is approximately
$\begin{split}I_{1}\approx\int_{0}^{a}2\pi r_{1}e^{-r_{1}^{2}\pi}e^{r_{1}^{2}(-\frac{\pi}{3}-\frac{\sqrt{3}}{2}+\epsilon(\frac{4\pi}{3}-\frac{9}{8})+o(\epsilon))}dr_{1}.\\ \approx\int_{0}^{a}2\pi r_{1}e^{-r_{1}^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}dr_{1}.\end{split}$ (19)
Using the substitution $u=r_{1}^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))$, we get that
$I_{1}\approx\frac{\pi}{(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big).$
With another application of a Taylor expansion, we can write $I_{1}$ in terms of $\epsilon$ as
$\begin{split}I_{1}\approx\frac{\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big)+\\ \epsilon\frac{\pi}{(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})^{2}}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big)+o(\epsilon).\end{split}$ (20)
Plugging the explicit formulas for $A_{1}$ and $A_{2}$, we get that $I_{2}$ simplifies to
$I_{2}=\int_{0}^{a}2\pi r_{1}e^{-r_{1}^{2}\pi}e^{-r_{1}^{2}(\frac{\pi}{3}+\frac{\sqrt{3}}{2})}dr_{1}.$
Applying the substitution $u=r_{1}^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})$, we end up with
$I_{2}=\frac{\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})}\Big).$
Combining the calculations for $I_{1}$ and $I_{2}$, we have that the probability $p_{o}$ is approximately
$\begin{split}p_{0}(a,\epsilon)\approx\frac{\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}\cdot e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})}\cdot\Big(1-e^{-a^{2}(\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big)\\ +\epsilon\frac{\pi}{(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})^{2}}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big)+o(\epsilon).\end{split}$ (21)
Applying the Taylor expansion to the exponential functions, we get
$\begin{split}p_{0}(a,\epsilon)\approx\frac{\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})}\Big(a^{2}(\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon)\Big)+\\ \epsilon\frac{\pi}{(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})^{2}}\Big(1-e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2}-\epsilon(\frac{4\pi}{3}+\frac{9}{8})+o(\epsilon))}\Big)+o(\epsilon)\end{split}$ (22)
Simplifying even more,
$p_{0}(a,\epsilon)\approx\epsilon\cdot\frac{a^{2}\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}\Big(e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})}+\frac{4\pi}{3}+\frac{17}{8}\Big)+o(\epsilon),$
where the second part of the product is a constant that depends on $a$. This implies that for each $a$, there exists $\epsilon$ small enough, which makes $p_{0}$ as small as needed.
Figure 11: The red square is closed if it or one of its 8 neighbors contains a point with a contraction ratio greater than $1-\epsilon$. Since the distance to the nearest neighbor is at most $a$, and the square has side length $2a$, starting from the red square we cannot cross the blue edge in 2 steps. The states of the red and the green square are independent of each other. By construction, the states of the squares and their ”shield” area need to be disjoint. Thus we have a 3-dependent setting.
Component 1:
Take a point $x$ from one of the open squares of side length $2a$. The state of a square depends on the points in the neighboring squares. Let us denote neighboring squares as layer 1 and the square in question, layer 0. Since we want layers 0 and 1 to be independent of other similar squares, we have a 3-dependent site percolation measure. This is illustrated in Figure 11, where the state of the red square is independent of the state of the green square.
In Lemma 2 we calculated the probability that the distance to the nearest point is less than $a$ and that the contraction ratio of the origin is greater than $1-\epsilon$. In Lemma 3, we calculate the probability $p_{c}$ of a closed square and show that it can be made arbitrarily small as $\epsilon\rightarrow 0$ and as $a\rightarrow\infty$. This will show that the open squares percolate for appropriate $a$ and $\epsilon$.
Lemma 3.
The probability $p_{c}$ of a square being closed can be made arbitrarily small for large enough contraction ratios and large enough $a$.
Proof.
A square is closed if it satisfies any of the following conditions:
  1. 1.
    It is empty of points, which happens with probability $e^{-\lambda 4a^{2}}=e^{-4a^{2}}$;
  2. 2.
    It contains a point whose nearest neighbor is at a distance greater than $a$. This happens with a probability less than
    $\mathbb{E}[\mbox{\# of points in the square}]\mathbb{P}^{0}[h(0,\Phi)>a]=$
    $4a^{2}\int_{a}^{\infty}2\pi re^{-r^{2}\pi}dr=4a^{2}e^{-4a^{2}\pi}.$
  3. 3.
    Finally it contains a point which has its nearest neighbor at distance less than $a$ but has a contraction ratio greater than $\rho$. This probability can be upper bounded by
[…]
    $\rho>1-\epsilon\mbox{ in a square of side 2a}]=$
    $4a^{2}\cdot\mathbb{P}^{0}[d(0,h(0,\Phi))<a\mbox{ and that the origin has }\rho>1-\epsilon]=$
    $4a^{2}\cdot p_{0}(a,\epsilon),$
where $p_{0}(a,\epsilon)$ was calculated in the Lemma 2.
Combining them together, we get that the probability that a square is closed, $p_{c}$, is less than
$p_{c}\leqslant e^{-4a^{2}}+4a^{2}e^{-a^{2}\pi}+a^{2}\cdot p_{0}(a)$
$p_{c}\leqslant 4e^{-a^{2}}+4a^{2}e^{-4a^{2}\pi}+a^{2}\cdot\Big(\epsilon\cdot\frac{a^{2}\pi}{\frac{4\pi}{3}+\frac{\sqrt{3}}{2}}\Big(e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})}+\frac{4\pi}{3}+\frac{17}{8}\Big)+o(\epsilon)\Big).$
If we take $\epsilon(a)=o\Big(a^{4}(e^{-a^{2}(\frac{4\pi}{3}+\frac{\sqrt{3}}{2})})\Big)$, then as $a\rightarrow\infty$, $p_{c}$ goes to 0.
Lemma 4.
[44, 15] There is a $p_{open}<1$ such that in any $k$-dependent site percolation measure on $\mathbb{Z}^{2}$ satisfying the additional condition that each edge is open with probability at least $p_{open}$, the probability that $|C_{0}|=\infty$ is positive.
We know that $p_{open}=1-p_{c}$, and since $p_{c}$ can be as small as we need, it follows that there exists $p_{open}$ such that the open squares percolate.
Component 2:
Figure 12: Figures from [29]. Left: A finite open cluster at the origin, surrounded by a closed circuit in the dual lattice. Right: An infinite path from the origin crosses an open square a.s.
Since closed squares do not percolate, there exists a closed circuit of open squares surrounding it [29]. Thus, there exists a closed circuit of open squares around the origin. An infinite forward path from 0 has to cross this circuit of open squares somewhere. This follows directly from the Konig’s lemma [39], which says that an infinite path has to to go infinity. Once it crosses an open square, by construction, all points in the forward set are at a distance less than $a$ from their nearest neighbor. An infinite path without cycles must hence cross infinitely many open squares.
Component 3:
In the next part of the proof, we need to study the relations between Palm probabilities corresponding to different random measures or point processes living on the same probability space and being jointly stationarity. We use mainly the Mass Transport Formula (Theorem 6.1.34. in [7]).
We split $\Phi$ into two disjoint sub-point processes, $\tilde{\Phi}$ and $\hat{\Phi}$, such that $\Phi=\tilde{\Phi}+\hat{\Phi}$. We call the points of $\tilde{\Phi}$ the ”good” points and the points in $\hat{\Phi}$ as the ”bad” points.
Let $\tilde{\Phi}$ be the sub-point process of $\Phi$ with points with
  1. 1.
    a nearest neighbor at distance at most $a$,
  2. 2.
    contraction ratio at most $\rho$.
By construction, open squares contain only ”good” points. Thanks to the percolation of open squares, all infinite paths starting from 0 contain infinitely many ”good” points.
Points that do not satisfy the criteria for $\tilde{\Phi}$ belong to $\hat{\Phi}$.
For the ”good” sub-point process, $\tilde{\Phi}$ we have the following lemma:
Lemma 5.
Under the Palm probability of $\tilde{\Phi}$, mean size of the Backward set of 0 under $\tilde{\Phi}$ is integrable a.s.
Proof.
Let us denote the Palm expectation of $\tilde{\Phi}$ by $\tilde{\mathbb{E}}^{0}$. For each point $x$ of $\Phi$, we send mass $d(h^{2}(y,\Phi),h(y,\Phi))$ to point $y$ if $y$ is in the forward set of $x$ and $y\in\tilde{\Phi}$. We will apply the mass transport formula with
$m(x,y,\omega)=d(h^{2}(y,\Phi),h(y,\Phi))\mathbf{1}\{y\in\tilde{\Phi},\mbox{and }y\in For(x,\Phi).\}.$
Then
$\lambda\mathbb{E}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\tilde{\Phi}(dy)\Big]=\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}d(h^{2}(y,\Phi),h(y,\Phi))\mathbf{1}\{y\in For(0,\Phi)\}\Phi(dy)\Big].$
Since
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}d(h^{2}(y,\Phi),h(y,\Phi))\mathbf{1}\{y\in For(0,\Phi)\}\Phi(dy)\Big]<a\sum_{i=0}^{\infty}(1-\epsilon)^{i}<\infty,$
it implies that
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(y,0,\omega)\Phi(dy)\Big]<\infty,$
which implies that the size of the backward set of the origin under $\tilde{\Phi}$ is integrable a.s. Furthermore,
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(y,0,\omega)\Phi(dy)\Big]\geqslant|Back(0,\tilde{\Phi})|\cdot a,$
which implies that the size of the backward set of the origin under $\tilde{\Phi}$ is finite a.s.
Now we’re ready for the final part of the proof.
We claim that the forward path has to be finite a.s. We proved that the number of open squares crossed by the forward path is bounded. Of course, a forward path crosses closed squares as well. So once the infinite path has crossed all its open squares, it can only cross closed squares. In addition, a path can then only go from one closed square to the neighboring closed square. However, the connected component of any closed square is a.s. finite. Therefore there does not exist a closed path that goes to infinity avoiding open squares. It follows that the number of closed squares in the forward path is also finite a.s.. We immediately get the desired result, namely the forward path is finite a.s.
Lemma 6.
The Backward set of 0 is a.s. finite under the Palm probability of $\Phi$.
Proof.
Let us denote the Palm expectation of $\hat{\Phi}$ as $\hat{\mathbb{E}}^{0}$. This proof again uses the mass transport principle in the following way:
$m(x,y,\omega)=\mathbf{1}\{x\in\hat{\Phi},\mbox{and }|Back(x,\hat{\Phi})|=\infty\}\mathbf{1}\{y\in\tilde{\Phi},y\in For(x,\Phi)\}.$
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\hat{\Phi}\}(dx)\Big]=$
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}\mathbf{1}\{|Back(x,\hat{\Phi})|=\infty\}\mathbf{1}\{y\in\tilde{\Phi},y\in For(x,\Phi)\}\hat{\Phi}(dx)\Big]$
$\hat{\lambda}\hat{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}\mathbf{1}\{x\in\hat{\Phi},\mbox{and }|Back(0,\hat{\Phi})|=\infty\}\mathbf{1}\{y\in For(0,\Phi)\tilde{\Phi}(dy)\Big].$
Which is the same as
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\hat{\Phi}(dy)\Big]=\hat{\lambda}\hat{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(x,0,\omega)\tilde{\Phi}(dx)\Big].$
Hence, the expected mass in, is the expectation under the Palm of $\tilde{\Phi}$ of the number of ”bad” points of the predecessors of 0 with an infinite backward set. As a consequence of Lemma 5, this expected mass in has to be 0. Hence, the expected mass out is also zero. The Palm probability of $\hat{\Phi}$ that the origin has a finite backward set a.s.
Here is an alternative proof of Lemma 6 based on the Point-Map cardinality classification [8]. Since the forward path is finite a.s., Follower party has to be $\mathcal{F}/\mathcal{F}$. Hence, the backward path has to be finite a.s.
Therefore we obtained the desired result and completed the alternative proof that the Poisson descending chain is finite a.s.
6.2 Follower Parties at Step 1
We now extend the proof to the first step of the dynamics.
Theorem 3.
The Poisson Follower Model at step 1 does not percolate.
Proof.
The idea of the proof is the same as in the proof of lack of percolation at step 0. At first glance, the proof for step 1 might seem repetitive to the reader. We want to stress that a first essential difference comes in the percolation argument. Since at step 1, the point process is not Poisson, we need to adjust the percolation argument. In addition, it is quite cumbersome to write out all the movements of points and get the probability for some contraction ratio of points of $\Phi_{1}$ as we did for step 0. Instead, and this is the second important difference, we use a combinatorial argument to show that the probability of having a large enough contraction ratio goes to 0 as the ratio goes to $1$. This argument is based on the fact that the factorial moment measure of order 3 of $\Phi_{1}$ has a bounded density. We use this to show that our new square tessellation percolates. Then, we use again Peierl’s argument to show that under the Palm probability of $\Phi_{1}$, the forward path from the origin has to hit an open square. In this case we have to add extra conditions for a square to be open. We show that the 4-dependent open boxes percolate. Therefore, if the forward set is infinite, it has to cross infinitely many open squares. Finally, we use the mass transport formula to show that the forward and the backward sets of the typical point are a.s.finite.
The contraction ratio is naturally extended to $\Phi_{1}$,
$\rho(x)=\frac{d(h(x,\Phi_{1}),h^{2}(x,\Phi_{1}))}{d(x,h(x,\Phi_{1}))}1_{h^{2}(x,\Phi_{1})\neq x}.$ (23)
For mutually closest neighbors in $\Phi_{1}$, we again define the contraction ratio to be 0.
Theorem 4.
The forward and the backward sets of 0 in $\Phi_{1}$ under $\mathbb{P}_{\Phi_{1}}^{0}$ are a.s. finite.
Proof.
The proof is based on a series of lemmas.
For the square lattice tessellation, we again use a square of side length $2a$. Let $\epsilon>0$ be the parameter in the contraction ratio.
In this proof we need to look at more than just the square and its neighbors. In order to make the discussion easier, we will call the square in focus, layer 0, the first 8 nearest squares, layer 1, the next 16 nearest squares, layer 2, etc. The different layers are illustrated in Figure 13 in different colors. For example, layer 1, which has 8 squares is in yellow.
Figure 13: Square and its layers. Layer 0 is shown in red, layer 1 is in yellow, layer 2 is in green and layer 3 is in gray
The definition of open square differs from that at step 0:
We want the open squares under $\Phi_{1}$ to have the same properties as the open squares at step 0 under $\Phi_{0}$. That is, the conditions on $\Phi_{1}$ are:
  1. (i)
    The square of layer 0 contains at least one point of $\Phi_{1}$, and each square in layer 1 contains at least one point of $\Phi_{1}$.
  2. (ii)
    Each point of $\Phi_{1}$ in the square of layer 0 has its nearest neighbor at distance at most $a$, where $2a$ is the side length of the squares. We have the same requirement for points in layer 1.
  3. (iii)
    Each point in the squares in layers 0 and 1 has a contraction ratio of at most $\rho$ with respect to $\Phi_{1}$.
These conditions are meant to provide the desired properties, as above, namely any follower path of $\Phi_{1}$ has to cross open squares and when it does, the path has steps of at most $a$.
In order to have $k$-dependency, we add the following conditions on points of $\Phi_{0}$:
  1. (a)
    For each square in layers 0 and 1, add an inscribed square in the center with side length $a$. Each of the inscribed squares should have at least one point of $\Phi_{0}$. In addition, each point of $\Phi_{0}$ in the squares of layers 0 and 1 has a nearest neighbor at a distance at most $a$. See Figure 14 for an illustration.
  2. (b)
    Each square in layers 2 and 3 has at least one point of $\Phi_{0}$ and all points of $\Phi_{0}$ in these squares have a nearest neighbor at a distance at most $a$.
    Figure 14: Additional conditions on $\Phi_{0}$ for layer 0 and layer 1. Each inner square in the layer 0 and 1 has at least one point.
Condition $(a)$ on $\Phi_{0}$ ensures $(i)$ on $\Phi_{1}$. Since the distances to the nearest points under $\Phi_{0}$ in layers 040-4 are less than $a$, these distances will remain less than $a$ under $\Phi_{1}$ in layers 0 and 1 as well. So conditions $(a)$ and $(b)$ on $\Phi_{0}$ ensure condition $(ii)$ on $\Phi_{1}$.
A square is closed if it not open. The proof of $4$-dependence is established in Lemma 7.
Then in Lemma 9, we show that the probability that a square is open can be made arbitrarily close to 1 by choosing $a$ and $\epsilon$ properly.
We can again break the proof of Theorem 4 into three separate components:
  1. 1.
    For a proper choice of lattice edge $2a$, there is a percolation of the open squares;
  2. 2.
    Each infinite path from 0 crosses infinitely many open squares;
  3. 3.
    Use Mass Transport Formula to show that the forward path from 0 has to be finite and then show that the backward set from 0 also has to be finite a.s.
Component 1: Percolation of open squares
This is the main difference with the proof of step 0. For step 0, we relied on the independence of the Poisson point process and could argue that there is 3-dependence. Now we show that the states of the squares as defined above are 9-dependent.
Let us call outer square the square that contains all the layers of elementary squares up to and including layer 4 and inner square the square which contains all elementary squares of layers 0 and 1.
Lemma 7.
Under the definitions for the open squares of this section, the site percolation model is 9-dependent.
Proof.
We want to show two things. Namely
  1. (j)
    The locations of points of $\Phi_{1}$ in layers 0 and 1 are a function of the points of $\Phi_{0}$ in layers 040-4.
  2. (j)
    The contraction ratios of points of $\Phi_{1}$ in layers 0 and 1 are a function of the points of $\Phi_{0}$ in layers 040-4.
To prove (j) it is enough to prove that no point of $\Phi_{0}$ outside the outer square penetrates the inner square of step 1.
Lets pick an arbitrary point of $\Phi_{0}$ which is outside the outer square and at distance $l$ from the outer square. Since every small square in layer 4 has a point of $\Phi_{0}$, the distance to the nearest point from layer 4, $d$, is less than $l+2a\sqrt{2}$. This follows directly from the triangle inequality. On the other hand, in order for the same point to enter the inner square, the point has to move at least $l+4a$. This means, it needs to follow a point which is at least $2l+8a$ away. Therefore we get the condition that
$l+2a\sqrt{2}>2l+8a,$
which is impossible.
To prove (jj) is it enough to prove that for all points $x$ of $\Phi_{1}$ of layers 0 and 1, $h(x,\Phi_{1})$ and $h^{2}(x,\Phi_{1})$ are determined by the locations of the points of $\Phi_{0}$ inside the outer square.
The idea is that the points of $\Phi_{0}$ in layers 2 and 3 act as a ”shield” that prevent points from outside the outer square to enter the inner square at step 1.
Take a point $x$ from layers 0 or 1 in $\Phi_{1}$. From the condition (ii) on $\Phi_{1}$ on open squares, we know that its nearest neighbor is at a distance of at most $a$. Therefore the distance between $x$ and $h^{2}(x,\Phi_{1})$ is less than $2a$. Since the sides of squares have length $2a$, it follows immediately that both $h(x,\Phi_{1})$ and $h^{2}(x,\Phi_{1})$ have to be within the layers 0-2. So it is enough to prove that the position of points in $\Phi_{1}$ in layers 020-2 is determined by the points of $\Phi_{0}$ in layers 030-3. From the argument above, no point of $\Phi_{0}$ from outside the outer square, could have entered the $2^{nd}$ layer in $\Phi_{1}$. Therefore all the points of $\Phi_{0}$ that play a role in the contraction ratio of $x$ with respect to $\Phi_{1}$ are within the outer square.
Therefore, because we want two squares together with their ”shield” layers to be disjoint, we need the squares to be at least distance 9 apart. Thus, we have a 9-dependent model.
The ”shield” idea is inspired in part by the Loop and Shield Conditions of [26]. We calculate the probability that the distance to the nearest point is less than $a$ and that the contraction ratio of the origin is greater than $1-\epsilon$. Then, we calculate the probability $p_{open}$ of an open square and show that it can be made arbitrarily close to 1 as $\epsilon\rightarrow 0$ and as $a\rightarrow\infty$. This will show that the open squares percolate.
Lemma 8.
The Palm probability for $\Phi_{1}$ that the distance to the nearest point is less than $a$ and that the contraction ratio of the origin is greater than $1-\epsilon$ tends to 0 as $\epsilon\rightarrow 0$.
Proof.
We need to show that as $\epsilon\rightarrow 0$ the probability that a contraction ratio of 0 is greater than $(1-\epsilon)$ goes to 0, i.e.,
$\mathbb{P}^{0}_{\Phi_{1}}\Big[(1-\epsilon)d(0,h(0,\Phi_{1}))<d(h(0,\Phi_{1}),h^{2}(0,\Phi_{1}))|\,d(0,h(0,\Phi_{1}))<a\Big]\rightarrow 0,\epsilon\rightarrow 0,$ (24)
where $\mathbb{P}^{0}_{\Phi_{1}}$ denotes the Palm probability under $\Phi_{1}$. This is enough to prove percolation of good boxes under $\Phi_{1}$. Let $A$ denote the event $(1-\epsilon)d(0,h(0,\Phi_{1}))<d(h(0,\Phi_{1}),h^{2}(0,\Phi_{1}))<d(0,h(0,\Phi_{1}))<a$.
In order to prove Equation (24), we use the fact that the factorial moment measure of order 3 of $\Phi_{1}$ has a bounded density. The argument is combinatorial in nature and we leave proof in the Appendix section LABEL:app:density-mcnp. Therefore,
$\mathbb{P}^{0}(A)=\mathbb{E}[\sum_{X\in B_{1}}\mathbf{1}_{\theta_{x}A}]$
$=\mathbb{E}[\sum_{X\in B_{1}}\sum_{Y\in\Phi_{1}}\mathbf{1}_{B(X\rightarrow Y)=1}\sum_{Z\in\Phi_{1}}\mathbf{1}_{B(Y\rightarrow Z)=1}\mathbf{1}_{d(X,Y)\leqslant a}\mathbf{1}_{d(Y,Z)\leqslant a}]$
$\leqslant\mathbb{E}\Big[\sum_{X\in B_{1}}\sum_{Y\in\Phi_{1}}\sum_{Z\in\Phi_{1}}\mathbf{1}_{d(X,Y)\leqslant a}\mathbf{1}_{d(Y,Z)\leqslant a}\mathbf{1}_{(1-\epsilon)d(X,Y)\leqslant d(Y,Z)\leqslant d(X,Z)}\Big]$
$=|B_{1}|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\gamma^{3}(0,y,z)\mathbf{1}_{|y|<a}\mathbf{1}_{|z-y|\leqslant a}\mathbf{1}_{(1-\epsilon)|y|<|z-y|<|y|}dydz\underset{\epsilon\rightarrow 0}{\rightarrow}0,$
where $\gamma^{3}$ denotes the $3^{rd}$ moment density of $\Phi_{1}$ and $B_{1}$ is a ball of volume 1. The relation holds since the volume of the region $\Big((1-\epsilon)|y|<|z-y|<|y|\Big)$ goes to 0 as $\epsilon\rightarrow 0$.
Let us denote by $p_{ct}(a)$, the probability that the origin has a contraction ratio less than $1-\epsilon$ under $\Phi_{1}^{0}$.
Now that we have a relation between $\epsilon$ and the probability for a point to have a contraction ratio greater than $1-\epsilon$, we can find the probability $p_{open}$ that the square is open.
Lemma 9.
The probability $p_{open}$ of a square being open can be made arbitrarily large for large enough contraction ratios and large enough $a$.
Proof.
Again, a square is open if it satisfies all of the following conditions:
  1. i
    For each square in layer 0 and 1, add an inscribed square in the center with an side length $a$. Each of the inscribed squares should have at least one point and a nearest neighbor at a distance at most $a$. The probability that the 9 center squares of side length $a$ have at least one point is
[…]
  2. $i$
    Each point in the square has its nearest neighbor at distance at most $a$, where $2a$ is the side length of the squares. We have the same requirement for the layers 0, 1, 2 and 3. The probability that one point has its nearest neighbor at a distance at most $a$ is
    $\int_{0}^{a}2\pi re^{-r^{2}\pi}dr=1-e^{-a^{2}\pi}.$
    In total we have 40 squares, which are all independent of each other in $\Phi_{0}$.
  3. iii
    Each point in the square and the neighboring 8 squares has a contraction ratio of at most $\rho$ in $\Phi_{1}$. This happens with probability greater than
[…]
    $36a^{2}\cdot\mathbb{P}^{0}[d(0,h(0,\Phi_{1}))<6a\mbox{, and that the origin has }\rho<1-\epsilon]=$
    $36a^{2}\cdot p_{ct}(a,\epsilon),$
    where $p_{ct}(a,\epsilon)$ is the probability that the origin has a contraction ratio of less than $1-\epsilon$ under $\Phi_{1}^{0}$.
Combining them together, we get that the probability that a square is open $p_{open}$, is greater than the product of the probabilities written above. As $\epsilon\rightarrow 0$ and $a\rightarrow\infty$, $p_{open}\rightarrow 1$.
Lemma 10.
[15, 44] There is a $p_{open}<1$ such that in any $k$-dependent site percolation measure on $\mathbb{Z}^{2}$ satisfying the additional condition that each edge is open with probability at least $p_{open}$, the probability that $|C_{0}|=\infty$ is positive.
We know from Lemma 9 that $p_{open}$, can be as close to $1$ as we need, therefore there exists $p_{open}$ such that it percolates. The proof of Lemma 10 can be found in [15].
Therefore, if we take $p_{open}$ large enough, it follows from $k$-dependent percolation that the set of open squares percolates.
Component 2:
Again, since closed squares do not percolate, there exists a closed circuit of open squares surrounding it [29]. Thus, there exists a closed circuit of open squares around the origin. An infinite forward path from 0 has to cross this circuit of open squares somewhere. Once it crosses an open square, by construction, all points in the forward set are at a distance less than $a$ from their nearest neighbor. An infinite path without cycles must hence cross infinitely many open squares.
Component 3:
This part of the proof is also the same as for step 0, just applied to $\Phi_{1}$. We again study the relations between Palm probabilities corresponding to different random measures or point processes living on the same probability space and being jointly stationarity. We use mainly the Mass Transport formula (Theorem 6.1.34. in [7]).
We split $\Phi_{1}$ into two disjoint sub-point processes, $\tilde{\Phi}_{1}$ and $\hat{\Phi}_{1}$, such that $\Phi_{1}=\tilde{\Phi}_{1}+\hat{\Phi}_{1}$. We call the points of $\tilde{\Phi}_{1}$ the ”good” points and the points in $\hat{\Phi}_{1}$ as the ”bad” points.
Let $\tilde{\Phi}_{1}$ be the sub-point process of $\Phi_{1}$ with points with
  1. 1.
    a nearest neighbor at distance at most $a$,
  2. 2.
    contraction ratio in $\Phi_{1}$ is at most $\rho$.
By construction, open squares contain only ”good” points. Thanks to the percolation of open squares, all infinite paths starting from 0 contain infinitely many ”good” points.
Points that do not satisfy the criteria for $\tilde{\Phi}_{1}$, belong to $\hat{\Phi}_{1}$.
For the ”good” sub-point process, $\tilde{\Phi}_{1}$ we have the following lemma:
Lemma 11.
Under the Palm probability of $\tilde{\Phi}_{1}$, the mean size of the Backward set of 0 under $\tilde{\Phi}_{1}$ is finite a.s.
Proof.
Let us denote the Palm expectation of $\tilde{\Phi}_{1}$ by $\tilde{\mathbb{E}}^{0}$. For each point $x$ of $\Phi_{1}$, we send mass $d(h^{2}(y,\Phi),h(y,\Phi))$ to point $y$ if $y$ is in the forward set of $x$ and $y\in\tilde{\Phi}_{1}$. We will apply the mass transport formula with
$m(x,y,\omega)=d(h^{2}(y,\Phi_{1}),h(y,\Phi_{1}))\mathbf{1}\{y\in\tilde{\Phi}_{1},\mbox{and }y\in For(x,\Phi_{1}).\}.$
Then
$\lambda\mathbb{E}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\tilde{\Phi_{1}}(dy)\Big]=\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}d(h^{2}(y,\Phi_{1}),h(y,\Phi_{1}))\mathbf{1}\{y\in For(0,\Phi_{1})\}\Phi_{1}(dy)\Big].$
Since
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}d(h^{2}(y,\Phi_{1}),h(y,\Phi_{1}))\mathbf{1}\{y\in For(0,\Phi_{1})\}\Phi_{1}(dy)\Big]<a\sum_{i=0}^{\infty}(1-\epsilon)^{i}<\infty,$
it implies that
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(y,0,\omega)\Phi_{1}(dy)\Big]<\infty,$
which means that the size of the backward set of the origin under $\tilde{\Phi}_{1}$ is integrable a.s.
Additionally, we again have that
$\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(y,0,\omega)\Phi_{1}(dy)\Big]\geqslant|Back(0,\tilde{\Phi}_{1})|\cdot a,$
which implies that the size of the backward set of the origin under $\tilde{\Phi}_{1}$ is finite a.s.
Finally, we’re ready for the final part of the proof. This part of the proof is the same as for Follower Model at step 0.
We again claim that the forward path has to be finite a.s. We proved that the number of open squares crossed by the forward path is bounded. Of course, a forward path crosses closed squares as well. So once the infinite path has crossed all its open squares, it can only cross closed squares. Additionally, a path can only cross from one closed square to the adjacent closed squares. However, the connected component of closed squares is finite. Therefore there does not exist a closed path that goes to infinity avoiding open squares. It follows that forward path through closed squares is also finite a.s. We immediately get the desired result, namely the forward path is finite a.s.
Lemma 12.
The Backward set of 0 is a.s. finite under the Palm of $\Phi_{1}$.
Proof.
Let us denote the Palm expectation of $\hat{\Phi}_{1}$ as $\hat{\mathbb{E}}^{0}$. This proof again uses the mass transport principle in the following way:
$m(x,y,\omega)=\mathbf{1}\{x\in\hat{\Phi}_{1},\mbox{and }|Back(x,\hat{\Phi_{1}})|=\infty\}\mathbf{1}\{y\in\tilde{\Phi_{1}},y\in For(x,\Phi_{1})\}.$
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\hat{\Phi_{1}}\}(dx)\Big]=$
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}\mathbf{1}\{|Back(x,\hat{\Phi}_{1})|=\infty\}\mathbf{1}\{y\in\tilde{\Phi}_{1},y\in For(x,\Phi_{1})\}\hat{\Phi}_{1}(dx)\Big]$
$\hat{\lambda}\hat{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}\mathbf{1}\{x\in\hat{\Phi}_{1},\mbox{and }|Back(0,\hat{\Phi}_{1})|=\infty\}\mathbf{1}\{y\in For(0,\Phi_{1})\tilde{\Phi}_{1}(dy)\Big].$
Which is the same as
$\tilde{\lambda}\tilde{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(0,y,\omega)\hat{\Phi}_{1}(dy)\Big]=\hat{\lambda}\hat{\mathbb{E}}^{0}\Big[\int_{\mathbb{R}^{2}}m(x,0,\omega)\tilde{\Phi}_{1}(dx)\Big].$
Hence, the expected mass in, is the expectation under the Palm of $\tilde{\Phi}_{1}$ of the number of ”bad” points of the predecessors of 0 with an infinite backward set. As a consequence of the Lemma 11, this expected mass in has to be 0. Hence, the expected mass out is also zero. The Palm probability of $\hat{\Phi}_{1}$ that the origin has a finite backward set a.s.
Again, we have an alternative proof of Lemma 12 based on the Point-Map cardinality classification [8]. Since the forward path is finite a.s., Follower party has to be $\mathcal{F}/\mathcal{F}$. Hence, the backward path has to be finite a.s. Which is what we wanted to show. ∎
Therefore, by extending the proof of step 0, we have shown that the Follower dynamics at step 1 does not percolate.
6.3 Follower Parties at later steps
We can in principle extend the proof of lack of percolation of follower parties at step 1 to later steps. Comparing the proofs for step 0 and step 1, we see that an essentially different part is the $k$-dependent percolation argument. We again tessellate the plane with the square grid of side length $a$ and give the same set of conditions for a square to be open. Then, for each step, we need to know from how far away would a point come to an open square. Depending on the step of the dynamics, we can potentially find a $k$ and then apply the $k$-dependence argument.
In addition, for each step we need to show that the $3^{rd}$ factorial moment has a bounded density. This can be achieved by a combinatorial argument. The rest of the proof would be the same as for step 0 as it doesn’t use Poisson point process and is a general argument.
7 Asymptotic Regime
This section focuses on the limiting behavior of party trees. We analyze stable trees in Subsection 7.1. Stable trees are special types of Follower Parties, which are trees for which the graph structure doesn’t change with steps. Such trees have a special structure. As we shall see they have no branching outside the root. We conjecture that all the Follower Parties eventually become stable trees.
Finally, in Subsection 7.2 we show some general long term relations. When the number of steps of the Follower Dynamics tends to $\infty$, there are two subprocesses of the initial point process being formed. One converges in total variation- it is the ultimate leader pair point process, and the other one converges weakly to its limit, the ultimate follower point process. From simulations, in dimension 2, the density of ultimate leader pairs is approximately 0.66. The rest of the points, the ultimate followers, converge weakly to one of the ultimate leader pairs.
7.1 Stable Tree
We call a party tree stable if the leader/followers relationship in this tree does not change over time under the dynamics. An example of a stable tree is shown in Figure 15.
Figure 15: Left:Tree at the initial time step. Right: Zoomed in situation after 20 time steps
For such stable parties, we are able to write a closed form trajectory of each agent in the tree based on the location of its ultimate leaders. We will do calculations of positions after the first time step, because after one step of the dynamics there is a fixed pair of points, and it is easier to derive positions with respect to a fixed point. We will look at one tree of followers/leaders in the following way. Let us denote the position of the ultimate leader pair as $a_{0}=a_{0}^{(1)}$. Superscript $(i)$ denotes the time step of the overall dynamics. The opinion of a follower of order 1 will be denoted by $a_{1}=a_{1}^{(1)}$. Now we take a follower of $a_{1}$ and denote its opinion as $a_{2}=a_{2}^{(1)}$, etc. The opinion of a follower of order n1n-1 will be denoted as $a_{n}=a_{n}^{(1)}$. The position of agent $n$ at time step $i$ will be denoted by $a_{n}^{(i)}$.
Recall that by our dynamics and because our tree is stable we have that
$a_{n}^{(i)}=\frac{a_{n}^{(i-1)}+a_{n-1}^{(i-1)}}{2}.$
Thus we can derive a recursive formula for the position of any agent at any time step that depends only on the positions of the leaders of various order of that agent.
Lemma 13.
Under the setting described above, the position of agent $a_{n}$ at time step $i$ is given by
$a_{n}^{(i)}=\begin{cases}\frac{1}{2^{i}}\big(\sum_{k=0}^{i}{i\choose k}a_{n-k}\big),&i<n,\\ \frac{1}{2^{i}}\big(\sum_{k=0}^{n-1}{i\choose k}a_{n-k}+\sum_{k=n}^{i}{i\choose k}a_{1}\big),&i\geqslant n.\par\end{cases}$
Proof.
Follows by induction.
Note that in the limit, $i\rightarrow\infty$, $\frac{1}{2^{i}}\big(\sum_{k=0}^{n-1}{i\choose k}a_{n-k}\big)\rightarrow 0$ and $\frac{1}{2^{i}}\big(\sum_{k=n}^{i}{i\choose k}a_{1}\big)\rightarrow a_{1}$.
Proposition 1.
If the tree is stable, it cannot have branching outside of the root.
Proof.
The proof is by contradiction. Let us take one branch of the tree and look for the first branching off, denoting the point as $a_{n}$. Then its two followers are denoted as $a_{n+1,1}$ and $a_{n+1,2}$. Notice that from the Lemma 13, the formula for the position of $a_{n+1,1}^{(i)}$, of agent $(n+1,1)$ and of $a_{n+1,2}^{(i)}$, of agent $(n+1,2)$ at time $i\geqslant n+1$ are
$a_{n+1,1}^{(i)}=\frac{1}{2^{i}}\Big(a_{n+1,1}+\sum_{k=1}^{n}{i\choose k}a_{n+1-k}+\sum_{k=n+1}^{i}{i\choose k}a_{1}\Big)$
$a_{n+1,2}^{(i)}=\frac{1}{2^{i}}\Big(a_{n+1,2}+\sum_{k=1}^{n}{i\choose k}a_{n+1-k}+\sum_{k=n+1}^{i}{i\choose k}a_{1}\Big).$
So the distance between them $d(a_{n+1,1}^{(i)},a_{n+1,2}^{(i)})$ is
$d(a_{n+1,1}^{(i)},a_{n+1,2}^{(i)})=\frac{1}{2^{i}}d(a_{n+1,1},a_{n+1,2}).$
Thus the distance between $a_{n+1,1}$ and $a_{n+1,2}$ halves every time step.
First, we prove the result in the special case when $n=2$ and then prove in general with the same idea but more calculations.
Position of $a_{2}$ changes with time step $i$ in the following way
$a_{2}^{(i)}=\frac{a_{2}+(2^{i}-1)a_{1}}{2^{i}},$
while $a_{3,1}$ and $a_{3,2}$ move as
$a_{3,1}^{(i)}=\frac{a_{3,1}+ia_{2}+(2^{i}-i-1)a_{1}}{2^{i}}$
$a_{3,2}^{(i)}=\frac{a_{3,2}+ia_{2}+(2^{i}-i-1)a_{1}}{2^{i}}$
We want to show that there exists $i$ for which $d(a_{3,1}^{(i)},a_{2}^{(i)})>d(a_{3,1}^{(i)},a_{3,2}^{(i)})$ or $d(a_{3,2}^{(i)},a_{2}^{(i)})>d(a_{3,1}^{(i)},a_{3,2}^{(i)})$.
Note that
$d(a_{3,1}^{(i)},a_{2}^{(i)})=\frac{|a_{3,1}+(i-1)a_{2}-a_{1}|}{2^{i}}.$
So the question is, whether there is an $i$ such that
$\frac{|a_{3,1}+(i-1)a_{2}-a_{1}|}{2^{i}}>\frac{d(a_{3,1},a_{3,2})}{2^{i}},$
or after multiplying by $2^{i}$
$|a_{3,1}+(i-1)a_{2}-a_{1}|>d(a_{3,1},a_{3,2}).$
We will denote the $x$-coordinates of $a_{n}$ as $x_{n}$ and $y$-coordinates as $y_{n}$ respectively. Thus
$|a_{3,1}+(i-1)a_{2}-a_{1}|=|(x_{3,1}+(i-1)x_{2}-x_{1},y_{3,1}+(i-1)y_{2}-y_{1})|.$
So after squaring the inequality, the question is, whether there is an $i$ such that
$(x_{3,1}-x_{1}+(i-1)x_{2})^{2}+(y_{3,1}-y_{1})+(i-1)y_{2})^{2}>d(y_{3},y_{3})^{2}.$
After rearranging the terms and grouping them together, we get an inequality of the form
$(i-1)^{2}(x_{2}^{2}+y_{2}^{2})+(i-1)b+c>0,$
where $b=2((x_{3,1}-x_{1})x_{2}-(y_{3,1}-y_{1})y_{2})$, and $c=(x_{3,1}-x_{1})^{2}+(y_{3,1}-y_{1})^{2}-d(x_{3},y_{3})^{2}$.
Since it is a degree 2 polynomial with positive square coefficient ($x_{2}^{2}+y_{2}^{2}>0$), it follows directly that it will be positive for some $i$, which is what we wanted to show.
Now in a similar way we extend this result to an arbitrary $n$. We want to show that there exists $i$ such that
$d(a_{n+1,1}^{(i)},a_{n}^{(i)})>d(a_{n+1,1}^{(i)},a_{n+1,2}^{(i)}).$
Recall that for $i>n$
$a_{n}^{(i)}=\frac{1}{2^{i}}\Big(\sum_{k=0}^{n-1}{i\choose k}a_{n-k}+\sum_{k=n}^{i}{i\choose k}a_{1}\Big),$
and
$a_{n+1,1}^{(i)}=\frac{1}{2^{i}}\Big(\sum_{k=0}^{n}{i\choose k}a_{n+1-k}+\sum_{k=n+1}^{i}{i\choose k}a_{1}\Big).$
From the initial formulas, we know that the distance changes as
$\begin{split}d(a_{n+1,1}^{(i)},a_{n}^{(i)})=\Big|\frac{(\sum_{k=0}^{n}{i\choose k}a_{n+1-k}+\sum_{k=n+1}^{i}{i\choose k}a_{1})}{2^{i}}-\frac{(\sum_{k=0}^{n-1}{i\choose k}a_{n-k}+\sum_{k=n}^{i}{i\choose k}a_{1})}{2^{i}}\Big|.\end{split}$ (25)
Looking at the left hand side we see the summation
$\begin{split}\Big|\frac{a_{n+1,1}+{i\choose 1}a_{n}+{i\choose 2}a_{n-1}+...+{i\choose n}a_{1}+{i\choose n+1}a_{1}+\ldots+{i\choose i}a_{1}-({i\choose 0}a_{n}+{i\choose 1}a_{n-1}+\ldots+{i\choose n-1}a_{1}+\ldots+{i\choose i}a_{1})}{2^{i}}\Big|.\end{split}$ (26)
Many terms with $a_{1}$ cancel and the equation can be written compactly as
$d(a_{n+1,1}^{(i)},a_{n}^{(i)})=\frac{\Big|\sum_{k=0}^{n-2}{i\choose k}(a_{n+1-k}-x_{n-k})-{i\choose n-1}a_{1}\Big|}{2^{i}}.$
The distance between $a_{n+1,1}^{(i)}$ and $a_{n+1,2}^{(i)}$ changes as
$d(a_{n+1,1}^{(i)},a_{n+1,2}^{(i)})=\frac{d(a_{n+1,1},a_{n+1,2})}{2^{i}}.$
Thus we want to show that there exists a time step $i$ such that
$\Big|{i\choose n-1}a_{1}-\sum_{k=0}^{n-2}{i\choose k}(a_{n+1-k}-a_{n-k})\Big|>d(a_{n+1,1},a_{n+1,2}).$
Each of the $i\choose k$} is a polynomial of $i$. What we want is to somehow select a highest order polynomial. Take $i$ to be at least twice as $n$, i.e., $i>2n$. In that case, the highest degree polynomial will be for $i\choose n-1$} and is a polynomial of degree n1n-1 with positive coefficient next to $i^{n-1}$ term. We will denote the polynomial ${i\choose n-1}a_{1}$ as $P_{n-1}(i)$, and $\sum_{k=0}^{n-2}{i\choose k}(a_{n+1-k}-a_{n-k})$ as $Q(i)$. Note that $Q(i)$ is a polynomial of order less than n1n-1.
Like in the case $n=2$ we will denote the $x$-coordinates as $P_{n-1,1}(i)$ and $Q_{1}(i)$, and the $y$-coordinates as $P_{n-1,2}(i)$ and $Q_{2}(i)$, respectively. So now the inequality we are trying to show can be written more compactly as
$|(P_{n-1,1}(i)+Q_{1}(i),P_{n-1,2}(i)+Q_{2}(i))|>d(a_{n+1,1},a_{n+1,2}).$
Again square everything to get
$(P_{n-1,1}(i)+Q_{1}(i))^{2}+(P_{n-1,2}(i)+Q_{2}(i))^{2}>d(a_{n+1,1},a_{n+1,2})^{2}.$
After multiplying and rearranging we have
$(P_{n-1,1}(i)^{2}+P_{n-1,2}(i)^{2})+R>0,$
where
$R=2(P_{n-1,1}(i)Q_{1}(i)+P_{n-1,2}(i)Q_{2}(i))+(Q_{1}(i)^{2}+Q_{2}(i)^{2}-d(a_{n+1,1},a_{n+1,2})^{2}).$
$R$ is a polynomial of order less than 2(n1)2(n-1) and $P_{n-1,1}(i)^{2}$ and $P_{n-1,2}(i)^{2}$ are both polynomials of order 2(n1)2(n-1). Thus we can factor the inequality
$i^{2(n-1)}\Big(\frac{P_{n-1,1}(i)^{2}+P_{n-1,2}(i)^{2})+R}{i^{2(n-1)}}\Big)>0,$
i.e.
$i^{2(n-1)}\Big(\frac{P_{n-1,1}(i)^{2}}{i^{2(n-1)}}+\frac{P_{n-1,2}(i)^{2}}{{i^{2(n-1)}}}+\frac{R}{{i^{2(n-1)}}}\Big)>0.$ (27)
Notice that in Equation (27), the first two terms tend to constants as $n\rightarrow\infty$, and the last term tends to 0. Therefore the equation on the left hand side is greater than 0, for $i$ large enough. Thus for every $n$ there exists an $i$ such that the distance between $a_{n+1,1}^{(i)}$ and $a_{n+1,2}^{(i)}$ becomes smaller than the distance between $a_{n+1,1}^{(i)}$ and $a_{n}^{(i)}$. Therefore, the chain has to break at some point if it has branching at some point other than the root.
We conjecture based on the Proposition 1 is that, in the long term, there cannot be branching outside of the root in any tree.
Conjecture 1.
For every ultimate leader pair, there exists a step $k$ after which the party tree of the ultimate leader pair is stable.
Conjecture 2.
Limiting objects are ultimate leader pairs and ultimate followers.
7.2 Long Term Relations
In this subsection we look at some mass transport relations that hold in general in the limiting process. Limiting objects consist of ultimate leader pairs and ultimate followers, we can derive a relation between them. Let $\lambda=1$ be the intensity of the initial Poisson point process. Let $\lambda_{f}$ be the intensity of the ultimate followers and $\lambda_{l}$ be the intensity of the ultimate leaders. Take $\bar{N}$ to be the mean number of ultimate followers of one ultimate leader. Then by the mass transport principle it directly follows that
$1\cdot\lambda_{f}=\bar{N}\lambda_{l},$
using that $\lambda_{f}+\lambda_{l}=1$, implies that
$\bar{N}=\frac{1}{\lambda_{l}}-1.$
Since there are two ultimate leaders in a pair, the total number of ultimate followers is $N=2\bar{N}$. This relation should hold regardless of the dimensions.
Now if we plug in $\lambda_{l}=\frac{2}{3}$, like obtained in the simulations, we get that $N=1$. Implying that the mean size of a party is 3 in the limit, which is also consistent with the simulations.
Conjecture 3.
There exist ultimate followers.
8 Conclusion
In summary, we introduced a new model inspired by problems in opinion dynamics. We described various phenomena related to this dynamics and examined the system’s long-term behavior. We provided a systematic way to calculate the frequencies of specific configurations, such as the densities of ultimate leader pairs of order zero. This method also allows us to compute the probability of an agent leaving its party. Additionally, we numerically analyzed the long-term behavior of parties and derived general results about the shapes of limiting follower parties. We also show that follower parties are finite in size at steps 0 and 1. For future work, we plan to prove that follower parties stay finite at all steps, supported by simulations. Furthermore, we aim to demonstrate that the limiting objects can only be either ultimate leader pairs or ultimate followers. We also plan to explain the observed ratio between the densities of the ultimate leader pair point process and the ultimate follower point process. Finally, we hope to extend these results to higher dimensions.