### v4

# On The Poisson Follower Model

Abstract
We introduce a stochastic geometry dynamics inspired by opinion dynamics that captures the essence of modern asymmetric social networks with leaders and followers. Points in the Euclidean space represent opinions, and the leader of an agent is the one with the closest opinion. In this dynamics, each follower updates its opinion by halving the distance to its leader. We demonstrate that this simple dynamics and its iterations exhibit several interesting purely geometric phenomena related to the evolution of leadership and opinion clusters, which resemble those observed in social networks. We also show that when the initial opinions are randomly distributed as a stationary Poisson point process, the spatial frequency of each of these phenomena can be expressed through an integral geometry formula involving semi-algebraic domains. Finally, we analyze numerically the limiting behavior of this follower dynamics. In the Poisson case, the agents fall into two categories: ultimate followers, who continue updating their opinions indefinitely, and ultimate leaders, who adopt a fixed opinion after a finite time. Spatial discrete event simulations support all our findings.
1 Introduction
People’s opinions and beliefs are influenced in complex ways by families, friends, colleagues, and media, as well as by politicians and other mega-influencers [1, 11, 12, 23, 17, 20]. In recent decades, attempts have been made to understand aspects of this process using mathematical modeling and computational simulation such as surveys on opinion dynamics [3, 2, 37, 40, 44, 45].
Many models of opinion dynamics are based on the assumption that we are influenced more easily by people with whom we almost agree than by those whose views starkly differ from ours. A similar but more general phenomenon, known as biased assimilation, is recognized among psychologists; referring to our tendency to filter and interpret information in a way that supports our preconceived notions [34]. Models of opinion dynamics built on this assumption are known as bounded confidence models [3, 19, 39]. A popular example is the Hegselmann and Krause model [25, 26], which builds on earlier work by Krause [29, 30]. It has been extensively studied in the literature [35, 36, 14, 46], and will serve as our starting point here. The Hegselmann-Krause model examines the stochastic time evolution of a voter’s opinion in response to the opinion of other like-minded voters.
The original Hegselmann-Krause model is discrete in both time and opinion space with a finite number of agents moving one step at a time. Similar models that are continuous in time [43], opinion space [51], or opinion space and time [16, 22] have been proposed since then.
This paper proposes a mathematical model allowing one to analyze both short and long term opinion dynamics of a countable collection of agents. At any time step of the dynamics, each agent has an $\mathbb{R}^{d}$-valued opinion and can only be influenced by the agent whose opinion is closest to its opinion, which is referred to as its leader. This new model is meant to capture the essence of such social network interactions as Instagram and Twitter, where the asymmetric leader/follower dynamics is central.
The follower/leader connections form a directed graph, where the direction goes from the follower towards the leader. Two agents which follow each other will be called a leader pair. This structure defines a directed graph with different connected components. We will also show that, in this Poisson model, there are no cycles longer than 2 a.s. We think of a group of agents that are part of the same connected component of this follower graph as a party.
If the initial opinions form a random point process, the directed graph is generated over a random point set, and is hence a random graph. In several parts of this work, we will restrict ourselves to the case where the points of the initial configuration are sampled according to a stationary Poisson point process over $\mathbb{R}^{d}$. This choice is dictated by tractability reasons (other initial conditions will be considered in future steps). For instance, in this Poisson case, the follower/leader random graph is called the Poisson Nearest Neighbor graph and parties are called Poisson descending trees. They have been studied extensively [24, 33]. An example of such a tree is depicted in Figure 1 for $d=2$.
Figure 1: Example of a Poisson descending tree
The initial random graph evolves over time based on the prescribed dynamics. More precisely, given an initial point process, at the first time step of the dynamics, each agent’s opinion moves halfway to that of its leader, which will be shown to be uniquely defined in the Poisson model. This gives a second point process on which a new follower/leader directed graph is defined. The dynamics is then applied to this new graph, which defines a third point process, and so on. The Poisson Follower Point Process of order $n$ is the point process obtained at the $n^{th}$ iteration of this dynamics when starting from a Poisson Point Process. At the $n^{th}$ iteration, the agents or points are the vertices of a graph that will be referred to as the Poisson Follower graph of order $n$. As we will see, interestingly a given agent may follow different leaders at different times, i.e., the structure of the Poisson follower graph changes with $n$.
This dynamics can be seen defined in contrast to the Hegselmann-Krause dynamics [25]. The main difference is that opinion updates consist in averaging each opinion with that of its leader (in contrast to averaging each opinion over the set of opinions in a ball around it as in the Hegselmann-Krause dynamics). Also, the dynamics is extended to a particle system setting [32], namely to a stationary point process of $\mathbb{R}^{d}$ (in contrast to the Hegselmann-Krause dynamics which is only discussed on compact domains so far).
One of the surprising facts is that this simple dynamics, purely determined by the relative positions of agents, leads to a large collection of phenomena to which one can associate natural names in view of what is observed in social networks. Here are some instances of the phenomena observed in this dynamics in the Poisson case and that are discussed in detail in Section 3:
  • One agent can be a leader to several agents. The maximal number depends on the dimension. In dimension two, for all point configurations, a point can be a leader for up to 6 followers. This number is at most 5 a.s. in the Poisson case. This is related to the kissing number [18].
  • Each initial party has a.s. exactly one ”loop”, i.e., only one pair of mutually closest neighbors, or equivalently a leader pair.
  • If two points form a leader pair, then, after one step, their opinions merge and will never change.
  • At each step of the dynamics, an agent either keeps the same leader or follows another agent. Similarly, a leader at any step can keep its followers, gain, or lose some followers.
  • Surprisingly, two agents in a follower/leader pair can swap their roles in one step of the dynamics. This follower inversion is reminiscent of the mentor/mentored scenario where the direction of leadership changes as time evolves.
  • At any step of the dynamics, agents can be arranged in connected components called parties bellow; finite parties can be characterized by their leader pairs.
  • As a result of this dynamics, at each step, agents either stay in their party or switch positions within a party, or change party. Equivalently, under a step of the dynamics, a party can either stay the same, split, loose, or gain new agents.
The identification of these purely geometry based dynamical phenomena is complemented by analytical and simulation results on the Poisson model. The paper starts with Section 2 where we give the notation and the point process background used in the rest of the document. Section 3 presents the dynamical phenomena. It is one of the main contributions of the paper.
The main probabilistic question studied in the paper is about the spatial frequency of the phenomena described above at any given step. In Section 4, we show that in the Poisson case, the frequencies in question admit representations in terms of integral geometry formulas involving integration domains which are semi-algebraic sets of the Euclidean space and integrands which are determined by the volume of certain unions of balls. Section 5 is a complement to Section 4 which discusses numerical methods to evaluate the integral geometry formulas alluded to above and spatial simulation methods for the complementary estimation of the frequencies in question.
The next question discussed in the paper is of dynamical system nature and concerns the limiting behavior of the Poisson follower point process of order $n$ and of the Poisson follower graph of order $n$, as $n$ tends to infinity. These questions are discussed in Section 6. We conjecture a classification of the points of the initial Poisson point process into two categories: ultimate followers, which remain followers forever and whose opinions only converge weakly to their limiting value, and utimate leaders whose opinion converges in total variation to their limiting value, which is hence reached in a finite number of steps. In the same section, we study stable trees, i.e., trees that do not change their structure with time.
2 The Follower Model and its Background
2.1 Overview of Point Process
Let $||\bullet||$ denote the Euclidean norm on $\mathbb{R}^{2}$ and $B(A,r)$ the open ball of radius $r$ and center $A$.
Let $S$ be the $\mathbb{R}^{2}$ space equipped with its Borel $\sigma$- algebra $\mathcal{B}$ and Lebesgue measure $\nu$. A point process $\Phi$ on $S$ is a random locally finite subset of $\Lambda$. One can also view $\Phi$ as a random counting measure on $\Lambda$, having the form $\Phi=\sum_{i\in\mathbb{N}}\delta_{X_{i}}$, where $\{X_{i}\}_{i\in\mathbb{N}}$ is a countable collection of points in $S$ with no accumulation points.
A point process is called simple if $\Phi(\{x\})\leqslant 1$ for all $x\in S$. A point process whose distribution is invariant under translations is called stationary. The intensity measure of a point process $\Phi$ is the measure on $S$ defined by
$\lambda(B)=\mathbb{E}[\Phi(B)],\,\,B\in\mathcal{B}(\Lambda).$
Definition 1.
For any counting measure $\mu=\sum_{i\in\mathbb{N}}\delta_{x_{i}}$ and $n\in\mathbb{N}$, its $n$-th power in the sense of products of measures is
$\mu^{n}=\sum_{(i_{1},...,i_{n})\in\mathbb{N}^{n}}\delta_{(x_{i_{1}},...,x_{i_{n}})}.$
Define the $n$-th factorial moment measure as the following counting measure on $(\mathbb{R}^{2})^{n}$.
$\mu^{(n)}=\sum_{(i_{1}\neq...\neq i_{n})}\delta_{(x_{i_{1}},...,x_{i_{n}})}.$
Definition 2.
Moment measures. For a point process $\Phi$ on $\mathbb{R}^{2}$, let $\Phi^{n}$ be the $n$-th power of $\Phi$ and $\Phi^{(n)}$ be the $n$-th factorial moment of $\Phi$. We call $M_{\Phi^{n}}=E[\Phi^{n}(B)]$ the $n$-th moment measure (the first moment measure is the mean measure) of $\Phi$ and $M_{\Phi^{(n)}}=E[\Phi^{(n)}(B)]$ the $n$-th factorial moment measure.
Moment measures provide important average structural properties of the process, such as level of clustering or repulsion.
Definition 3.
(Poisson point process) Let $\Lambda$ be a locally finite measure on l.c.s.h. space $\mathbb{G}$. A point process $\Phi$ is said to be Poisson with intensity measure $\Lambda$ if for all pairwise disjoint sets $B_{1},...,B_{j}\in\mathcal{B}(\mathbb{G})$, the random variables $\Phi(B_{1}),...,\Phi(B_{j})$ are independent Poisson random variables with respective means $\Lambda(B_{1}),...,\Lambda(B_{j})$; i.e. $\forall m_{1},...,m_{j}\in\mathbb{N}$,
$\mathbb{P}(\Phi(B_{1})=m_{1},...,\Phi(B_{j})=m_{j})=\prod_{i=1}^{j}\frac{\Lambda(B_{i})^{m_{i}}}{m_{i}!}e^{-\Lambda(B_{i})}.$ (1)
Definition 4.
(Homogeneous Poisson point process on $\mathbb{R}^{2}$) If $\Phi$ is a Poisson point process on $\mathbb{R}^{2}$ with intensity measure $\Lambda(dx)=\lambda\times dx$ where $\lambda\in\mathbb{R}^{*}_{+}$ and $dx$ denotes the Lebesgue measure, then $\Phi$ is called a homogeneous Poisson point process of intensity $\lambda$.
Throughout the paper, we denote by $\Phi=\sum_{i\in\mathbb{N}}\delta_{x_{i}}$ the homogeneous Poisson point process on $\mathbb{R}^{2}$ with intensity $\lambda$, which serves as initial condition to the dynamics. For all $n\geqslant 0$, the Poisson Follower point process of order $n$ will be denoted by $\Phi_{n}$. Note that $\Phi=\Phi_{0}$. Let $\mathbb{R}^{*}=\mathbb{R}\setminus\{0\}$.
For two points $x,y\in\Phi$, we denote by $B(x\rightarrow y)$ the open ball with center $x$ and radius $d(x,y)$, which is the distance between $x$ and $y$. To indicate the distance between two points $x$ and $y$ at the $n^{th}$ iteration, we will write $d(x,y,\Phi_{n})$.
Palm Theory
Informally, the Palm measures of a point process $\Phi$ at a point $x\in\Lambda$ is the probability measure of $\Phi$ conditioned on having a point at location $x$. For a more detailed discussion on the matter, see [7, 27].
2.2 Notation for the Follower Dynamics
Consider a counting measure on $\mathbb{R}^{2}$. Take two points $A$ and $B$ in this counting measure. If the closest point to $A$ is $B$, we say that $A$ follows $B$. Other equivalent statements are, ”$A$ is a follower of BB”, or ”$B$ is a leader of AA”, and ”$B$ is followed by AA”, etc. We will also use notation $A\rightarrow B$ to mean that $A$ follows $B$. If $B$ follows $C$, and $A$ follows $B$, we call $C$ a leader of order two of $A$. We call leader pair a pair of agents that follow each other. In a homogeneous Poisson point process, each point A has a.s. a unique leader and there are a.s. no other cycles involved than leader pairs [24].
Because of scale invariance of the Poisson point process, we will assume without loss of generality that $\lambda=1$ in subsequent sections.
3 Phenomena Observed in the Follower Dynamics
In this section, we give some basic definitions to be used throughout the paper. We also identify and set the terminology for the phenomena of interest and give some illustrations.
3.1 Agent Events
Here is a list of situations that can happen to an agent when applying the dynamics at any step.
Leader swap of an agent
The ”out” edge of this agent is either kept (leader keep) or changed (leader swap) by the dynamics. Notice that these two situations are mutually exclusive since, by construction, each point gets only one leader at any given time. For the examples in Figure 2, Agent 3 swaps its leader from 1 to 2, as a result of the dynamics.
Figure 2: The Step 0 positions are shown in black. Step 1 positions are in red, with the new positions denoted with the prime ’. Note that 0 and 1 become 1’ in both images. Left: Example of a Follower Loss/Follower Gain. Right: Example of formation of a leader pair of order 1. Same convention as in the left image. In both cases 4 experiences a leader swap whereas 2 experiences a leader keep.
Follower gain, loss of keep of an agent
Concerning the set of ”in” edges of the agent, one can have a follower loss, a follower gain or a follower keep. The last case is that where the set set of followers is the same before and after the considered step of the dynamics. Notice that a loss and a gain can happen at the same time. In the left of Figure 2, Agent 1 lost its follower, Agent 3; Agent 2 gained a follower, namely Agent 3.
Leader pair
We call initial leader pairs, agents involved in a leader pair of order 0 (i.e. at step 0). New leader pairs, forming at step 1. will be called leader pairs of order 1, etc.
Leader pair formation by an agent
This situation is that where an agent which is not in a leader pair becomes part of a leader pair at the next step of the dynamics. An example of this is shown in the right Figure 2.
Figure 3: Example of follower inversion. Again 0 and 1 merge to become 1’. The positions at the step 1 are shown in red, with the new positions denoted with the prime, like for example $1^{\prime}$.
Figure 4: Example of party fission
Follower inversion for an agent
Another special case is that where, after one step, the leader of an agent becomes its follower. We call such a situation a follower inversion. We do not count here situations where the leader and follower become a leader pair at step 1 as a follower inversion. An example of follower inversion is shown in Figure 3.1. Agent 3 was initially following Agent 2, but at step 1, Agent 2 follows Agent 3, whereas Agents 3 and 4 form a leader pair.
3.2 Parties
Forward and Backward Sets
We denote by $\mbox{For}(x,\Phi)$ the leader set of all orders of $x$ in $\Phi$, and we will call it the forward set of $x$. $\mbox{Back}(x,\Phi)$ denotes the follower set of all orders of $x$ in $\Phi$, and we call it the backward set of $x$. For each agent, there is a forward and a backward set at each step of the dynamics. See Figure 5 for these sets at step 0.
Figure 5: Representation of the forward and backward sets of an agent. The forward set of $x$ is connected to $x$ with dark gray arrows while the backward set of $x$ is with light gray ones. The agents connected by black arrows are in the party of $x$ but neither in the backward set nor in the forward set of $x$.
Parties
Roughly, we call party any connected component of the follower graph at any step. Note that two points $x$ and $y$ belong to the same party iff there exists a point $z$ (not necessarily different from $x$ or $y$), such that both $x$ and $y$ belong to the backward set of $z$, i.e., $x,y\in\mbox{Back}(z,\Phi)$.
Parties at step $k$
are defined as connected components at step $k$. We conjecture that all parties are a.s. finite at all steps. At any step, finite parties are characterized by their leader pair. Since a leader pair cannot be dissolved, the number of parties can only increase with the number of steps. An example can be found in Figure 5. Since the initial agent distribution is Poisson at step 0, every party is almost surely finite at step 0 [24].
Party Gain/ Party Loss
A given party (once it exists) can have fluctuations in its cardinality. A party gain is an increase in the party’s cardinality at a given step; this requires that at least one agent of another party joins. A party loss is a decrease in its cardinality at a given step; this requires at least one agent of this party to leave.
Party fission
The situation in which a new leader pair is created within a party is called a party fission. The initial party survives (its leader pair cannot be dissolved). A new party is however formed, associated with this new leader pair. Figure 4 gives an example of such a fission.
Party swap of an agent
A given agent can belong to different parties (as defined above) at different steps. For a given agent, a party swap is a change of party (namely a change of leader pair) at a given step. An example of party swap is shown in right of Figure 6, where agent $z$ and its followers are subject to a party swap. Note that an agent may experience a party swap even if it keeps its leader, but this leader is itself subject to a party swap (as for agent $x$ in this example). Another example of party swap of an agent is shown in left of Figure  6 where Agent A is subject to a party swap.
Party restructuring
By party restructuring, we mean a situation where the dynamics leads to a change within the directed graph describing the party, but to no change in the set of agents contained in the party, like the situation illustrated in left of Figure 2.
Party stability
A party is stable if when applying the dynamics, the set of agents involved is unchanged and the structure of the directed graph describing the party is unchanged.
Indefinitely stable party
A party is indefinitely stable if it is stable at any future time step. Stable parties exist (e.g. any linear directed path ending with a leader pair) and are discussed in Section 6.1.
Figure 6: Left: Example of a 4 body swap. A’,B’, $A_{1}^{\prime}$ and $B_{1}^{\prime}$ are the positions of agents in the next time step. Positions in the second time step are denoted by AA" and $A_{1}$" and connections are shown in green. Right: Example of a party swap. Initial positions are shown in black and step 1 is in red. Agent $z$ swaps leaders and parties. Together with $z$, agents $x$ and $v$ also experience a party swap.
4 body party swap
A 4 body party swap is an agent party swap that involves 4 different agents: $A,B,A_{1},$ and $B_{1}$, with, at step $k$, AA following $B$ and belonging to a party, and $A_{1}$ following $B_{1}$ and belonging to another party. Then, at step $k+1$, AA follows $A_{1}$ which follows $B_{1}$, and neither $B_{1}$ nor $B$ swap party. Hence $A$ is subject to a party swap. For the example of the left of Figure 6, for this party swap of $A$ to take place, the following inequalities must hold: $h^{2}<(\frac{h}{2}-l)^{2}+d^{2}$ (AA follows $B$ at step $k$), $d<\frac{h+j}{2}$ (AA follows $A_{1}$ at step $k+1$). It is easy to check that there are solutions to these inequalities. In this figure, $(B,C)$ forms a leader pair as well as $(B_{1},C_{1})$, so that this scenario is a party swap for $A$.
Figure 7: Step 0 positions and connections are shown in black. Step 1 connections are shown in red. Left: At step 0, $x$ follows $y$ whereas at step 1, $y$ follows $x$. We hence have a follower inversion. In addition, $x$ and $u$ become a leader pair of order 1. Right: Agent $v$ swaps parties and agents $u$ and $x$ form a leader pair of order 1.
Remark
Under the Follower Dynamics, in the Poisson case, we observe all the phenomena described in this section, and it is possible that after one step several events happen at the same time on a given set of agents. For example, at the agent level, in the left Figure 7, we observe both a follower inversion and the creation of a leader pair of order 1. At the agent level, in the right Figure 7, we observe both a party swap and the formation of a leader pair of order 1. The right of Figure 7 is also an example of a simultaneous party restructuring and party swap.
4 Frequency of the Phenomena
In this Section, we develop a systematic method to calculate or estimate exact values or bounds on the spatial frequencies of related events using integral geometry. We demonstrate that this involves evaluating integrals of specific exponential functions over semi-algebraic sets. Semi-algebraic sets are subsets of $\mathbb{R}^{2}$, which are finite unions of solutions to polynomial equations and inequalities with coefficients in $\mathbb{R}$ [31]. This approach applies to each of the phenomena listed above.
4.1 Integral Geometry Estimates of Densities of Leader Pairs
In this subsection, we give an integral geometry representation of leader pairs and then give numerical estimates. We illustrate the method of calculating the densities by showing a step by step calculation of the density of leader pairs in the initial configuration, i.e., the density of order 0. One can calculate frequencies of other configurations in the same manner. Then, we calculate an upper bound and an exact integral geometry formula for the density of leaders of order 1, type 1 or density of order one type 1. This is a known result [33], but we derive it here in a new way that can be extended to other phenomena. We also show how to calculate the density of leader pairs of order one. Again, by leader pairs, we mean a nearest neighbor pair.
4.1.1 Order Zero
Consider the point process $N^{(2)}$ consisting of pairs of points of $\Phi$ which are mutually closest points. Namely
$N^{(2)}=\sum_{i\neq j\in\mathbb{Z}}\delta_{x_{i},x_{j}}1_{\Phi(B(x_{i}\rightarrow x_{j}))=1},\\ 1_{\Phi(B(x_{j}\rightarrow x_{i}))=1},$
where $\Phi(B(x_{j}\rightarrow x_{i}))$ denotes the number of points in the open ball centered at $x_{j}$ with the radius $d(x_{j},x_{i})$.
The mean measure of $N^{(2)}$ is given by, for $A\subset\mathbb{R}^{2}\times\mathbb{R}^{2}$,
$\begin{split}\mathbb{E}[N^{(2)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y\in A}1_{x\neq y}1_{\Phi(B(x\rightarrow y))=1}1_{\Phi(B(y\rightarrow x))=1}\Phi^{(2)}({\rm d}x\times{\rm d}y)],\end{split}$ (2)
where $\Phi^{(2)}$ is the Poisson factorial moment measure of order 2 ([7] Section 3.3.2). By the higher order Campbell-Little-Mecke formula [5],
$\begin{split}\mathbb{E}[N^{(2)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y\in A}\mathbb{E}^{x,y}[1_{\Phi(B(x\rightarrow y))=1}1_{\Phi(B(y\rightarrow x))=1}]\lambda^{(2)}({\rm d}x{\rm d}y),\end{split}$ (3)
where $\mathbb{E}^{x,y}$ is the two point Palm expectation and $\lambda^{(2)}$ is the factorial Poisson moment measure of order 2 ([7] Section 3.3.2). Recall that for a Poisson point process, the $n^{th}$ factorial moment measure equals the $n^{th}$ power of the intensity measure. We will be using this over and over again in the calculations that follow.
Now by Slivnyak’s theorem [5]
$\mathbb{E}[N^{(2)}(A)]=\int 1_{x,y\in A}\mathbb{E}[1_{\Phi(B(x\to y))=0}\,1_{\Phi(B(y\to x))=0}]\lambda^{2}dxdy$ (4)
where $\Phi^{\prime}=\Phi+\delta_{x}+\delta_{y}$. Using the change of variables $(x,y)\rightarrow(x,u)$ with u=yxu=y-x and taking $A=C\times\mathbb{R}^{2}$, with $C\in\mathcal{B}(\mathbb{R}^{2})$, we get
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}1_{x\in C}1_{u\in\mathbb{R}^{2}}\mathbb{E}[1_{\Phi^{\prime}\circ\theta_{x}(B(o\rightarrow u))=1}1_{\Phi^{\prime}\circ\theta_{x}(B(u\rightarrow 0))=1}]\\ \lambda^{2}{\rm d}x{\rm d}u.$ (5)
The expectation is the probability for point $u$ to be the closest to 0 and for 0 to be closest to $u$. Then
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{\mathbb{R}^{2}}1_{u\in\mathbb{R}^{2}}\mathbb{E}[1_{\Phi(B(o\rightarrow u))=0}1_{\Phi(B(u\rightarrow 0))=0}]\\ \lambda^{2}{\rm d}u.$ (6)
Because of symmetry, and after switching to polar coordinates, we get:
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{\theta\in[0,2\pi]}\int_{0}^{\infty}ve^{-\lambda v^{2}\pi}e^{-\lambda v^{2}(\pi/3+\sqrt{3}/2)}\lambda^{2}{\rm d}v.$ (7)
Hence
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})]=|C|\int_{0}^{\infty}2\pi ve^{-\lambda v^{2}\pi}e^{-\lambda v^{2}(\pi/3+\sqrt{3}/2)}\lambda^{2}{\rm d}v.$ (8)
To evaluate the integral, we use the change of variables $v\rightarrow r$ with $r=\pi v^{2}$, and recall that $\lambda=1$. Then we get
$\mathbb{E}[N^{(2)}(C\times\mathbb{R}^{2})])=|C|\int_{0}^{\infty}e^{-r(\frac{\pi+(\pi/3+\sqrt{3}/2)}{\pi})}\rm dr=|C|\cdot\frac{\pi}{\pi+\pi/3+\sqrt{3}/2}\approx 0.62|C|.$
So the density of the ultimate leader pairs of order 0, i.e. points that after one step stay together forever, is $0.62$. Note that we are double counting the points since they are two points at one spot.
4.1.2 Density of Leader Pairs of Order One
In this subsection, we continue with similar calculations to determine the intensity of certain leader pairs of order $1$. First, we provide an upper bound on this density and then explain how to find the exact density. Each calculation requires the positions of at least $4$ points to know the positions at step 1, see Figure 8. The first situation involves two followers of one point that become a leader pair at step 1. An instance is given on the left of Figure 8. The second involves cases where a leader and a follower become a leader pair at the next step. An instance can be found on the right of Figure 8. Densities of both types are computed similarly, so we focus on the detailed calculation of type 1 only.
Figure 8: Left: Four points and their positions at step 1. Example of leader pair of order 1, type 1: two followers of one point become a leader pair at step 1. Right: Example of leader pair of order 1, type 2. A point and its leader become a leader pair at step 1.
Type 1
As mentioned above, leader pairs of order 1, type 1 are formed from the configurations where two step 0 followers of one agent become a leader pair at step 1. We first give an upper bound on the frequency of type 1 and then we analyze the exact frequency.
Let $\mathcal{D}\in{\mathcal{B}}^{4}$ be the set of all distinct points $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ that satisfy the conditions (9)-(12) defined below:
$d(x_{1},x_{3})<d(x_{1},x_{2}),\,d(x_{1},x_{3})<d(x_{1},x_{4}),$ (9)
which are the conditions needed for $x_{1}$ to follow $x_{3}$, in the absence of other points than these four points;
$d(x_{2},x_{3})<d(x_{2},x_{1}),\,d(x_{2},x_{3})<d(x_{2},x_{4}),$ (10)
which are the conditions needed for $x_{2}$ to follow $x_{3}$ as well, also in the absence of other points. Finally
$d(x_{3},x_{4})<d(x_{3},x_{1}),\,d(x_{3},x_{4})<d(x_{3},x_{2}),$ (11)
which are the conditions needed for $x_{3}$ to follow $x_{4}$ under the same conditions. The phenomenon we are interested in is that where, in the above configuration,
$d(x_{1}^{\prime},x_{2}^{\prime})<d(x_{1}^{\prime},x_{3}^{\prime}),\,d(x_{2}^{\prime},x_{1}^{\prime})<d(x_{2}^{\prime},x_{3}^{\prime}),$ (12)
where $x_{1}^{\prime}=\frac{x_{1}+x_{3}}{2}$, $x_{2}^{\prime}=\frac{x_{2}+x_{3}}{2}$, and $x_{3}^{\prime}=\frac{x_{3}+x_{4}}{2}$. That is $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are mutual closest neighbors in the absence of other points. Note that $\mathcal{D}$ is a semi-algebraic set [31].
For the points $(x_{1},...,x_{4})$ to be involved in the formation of a leader pair of type 1, it is necessary but not sufficient that $(x_{1},\ldots,x_{4})$ belong to $\cal D$}. For this to happen, in addition, it must be that there are no other points of the Poisson P.P. that change the facts that both $x_{1}$ and $x_{2}$ follow $x_{3}$, $x_{3}$ follows $x_{4}$, and $x_{1}^{\prime}$ and $x_{2}^{\prime}$ are mutually nearest neighbors.
Let $\Phi_{1,2,3,4}=\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ denote the point process $\Phi$ restricted to the set $x_{1},x_{2},x_{3}$, and $x_{4}$. Let $N_{1}^{(4)}$ be the point process of quadruples of points of the Poisson P.P. $\Phi$ that belong to $\cal D$}, and are such that the following event $M_{1}$ holds: in $\Phi$, both $x_{1}$ and $x_{2}$ follow $x_{3}$ and $x_{3}$ follows $x_{4}$.
In a first step, we evaluate the spatial frequency $\beta^{1}$ of the event $M_{1}$, which is an upper bound on the frequency of leader pairs of order 1, type 1.
For $A\in\mathcal{B}^{4}$, the mean measure of $N_{1}^{(4)}$ is given by,
$\mathbb{E}[N_{1}^{(4)}(A)]=\mathbb{E}[\sum_{x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}],$
where $\Phi$ is the P.P.P. and $B(x\rightarrow y)$ is the open ball of center $x$ and radius |xy||x-y|. In integral form, this is
$\mathbb{E}[N_{1}^{(4)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}$
$1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}\Phi^{(4)}(dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].$
By the higher order Campbell-Little-Mecke formula,
$\mathbb{E}[N_{1}^{(4)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}$
$\mathbb{E}^{x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}]\lambda^{(4)}(dx_{1}dx_{2}dx_{3}dx_{4}),$
where $\mathbb{E}^{x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x_{1},...,x_{4}$, and $\lambda^{(4)}$ is the Poisson factorial moment measure of order 4. By Slivnyak’s theorem,
$\begin{split}\mathbb{E}[N^{(4)}_{1}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x_{1},x_{2},x_{3},x_{4}\in A\cap\mathcal{D}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}]\lambda^{4}dx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
where $\hat{\Phi}=\Phi+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P of intensity $\lambda$. We can simplify the expression a bit more using stationarity of the Poisson point process. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C\in\mathcal{B}(\mathbb{R}^{2})$ a compact. Because we have a stationary point process,using the change of variables, $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
$\mathbb{E}[N_{1}^{(4)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{\cal D}}e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0))}\lambda^{4}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4}=|C|\beta_{1},$
with $\tilde{{\cal D}}=\{(x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}\}$.
This is the announced integral over a semi-algebraic set. Note that this integral is twice the density of interest since we do not distinguish whether $d(x_{1},x_{3})<d(x_{2},x_{3})$ or the other way around. In order to evaluate the integral, we need to write a formula for the volume of the union of 3 balls. This is discussed in Section 5.2.
Now that we have all the elements, we have a formula to calculate exact value of the volume $|\mathbb{U}(x_{1},x_{2},x_{3})|$ of the union of balls. For compactness, the integral of the frequency of type 1 can be written as:
$\mathbb{E}[N_{1}^{(4)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\lambda^{4}1_{\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in{\cal D}_{0}}e^{-\lambda|\mathbb{U}(x_{1},x_{2},x_{4})|}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4}:=|C|\beta^{1},$
with ${\cal D}_{0}={\cal D}\cap\{x_{3}=0\}$.
We now discuss the exact calculation. In order to get the event of interest, one should in addition have $\Phi_{1}(B(x_{1}\to x_{2}))=1$ and $\Phi_{1}(B(x_{2}\to x_{1}))=1$, with $\Phi_{1}$ the point process at step 1. In other words, certain refinements of the last configuration should be removed from the counting. We can order these refinements in the disjoint and exhaustive categories listed below
  1. 1.
    There is an extra point $x$ in $\Phi$ that follows $x_{1}$ in $\Phi$ and such that the distance from $x^{\prime}$ to $x_{1}^{\prime}$ is less than that from $x_{1}^{\prime}$ to $x^{\prime}_{2}$. The case where there is an extra point $x$ in $\Phi$ that follows $x_{2}$ in $\Phi$ and such that the distance $d(x,x_{2},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$ is analogous and is counted here as well. This is due to a fact that we do not distinguish whether $d(x_{1},x_{3})<d(x_{2},x_{3})$ or vice versa.
    For evaluating the frequency of this event, we have to enrich $\cal D$} by adding $x$ and state that $x$ follows $x_{1}$ and that none of the points $x_{1},x_{2},x_{3},x_{4}$ follows $x$, that is
    $\begin{split}d(x,x_{1})<d(x,x_{2}),\,d(x,x_{1})<d(x,x_{3}),\,d(x,x_{1})<d(x,x_{4}),\\ d(x_{1},x_{3})<d(x_{1},x),\,d(x_{2},x_{3})<d(x_{2},x),\,d(x_{4},x_{3})<d(x_{4},x),d(x_{3},x_{4})<d(x_{3},x).\end{split}$ (13)
    This adds 7 quadratic inequalities. Finally, the condition that $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$, gives one more quadratic inequality
    $d(x^{\prime},x_{1}^{\prime})<d(x^{\prime}_{1},x^{\prime}_{2}),$ (14)
    where $x^{\prime}=\frac{x+x_{1}}{2}$. So the frequency of this refinement can also be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{1}$, which is a refinement of $\cal D$} with one more variable and 8 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{5}$, the mean measure of $N_{1,1}^{(5)}$, which is the point process of the 5-tuples of points satisfying the above conditions, is given by,
    $\mathbb{E}[N_{1,1}^{(5)}(A)]$ $=$ $\mathbb{E}[\sum_{x,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}],$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,1}^{(5)}(A)]=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}\\ 1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}\Phi^{(5)}(dx\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    By the higher order Campbell-Little-Mecke formula,
    $\mathbb{E}[N_{1,1}^{(5)}(A)]$ $=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}$
    $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{1}))=1}]\lambda^{(5)}(dxdx_{1}dx_{2}dx_{3}dx_{4}),$
    where $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x,x_{1},...x_{4}$, and $\lambda^{(5)}$ is the factorial Poisson moment measure of order 5. By Slivnyak’s theorem,
    $\begin{split}\mathbb{E}[N_{1,1}^{(5)}(A)]=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{1}}\mathbb{E}[1_{\hat{\Phi}(B(x_{1}\rightarrow x_{3}))=0}\\ 1_{\hat{\Phi}(B(x_{2}\rightarrow x_{3}))=0}1_{\hat{\Phi}(B(x_{3}\rightarrow x_{4}))=0}1_{\hat{\Phi}(B(x\rightarrow x_{1}))=1}]\lambda^{5}dxdx_{1}dx_{2}dx_{3}dx_{4},\end{split}$
    where $\hat{\Phi}=\Phi+\delta_{x}+\delta_{x_{1}}+\delta_{x_{2}}+\delta_{x_{3}}+\delta_{x_{4}}$ is a Poisson P.P with intensity $\lambda$. Take $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C$ a compact. Because we have a stationary point process, using the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,1}^{(5)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{1}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow\tilde{x}_{1})))}\lambda^{5}d\tilde{x}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with $\tilde{{\cal D}_{1}}=\{(x-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x,\,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{1}\}$.
    Similar to the derivation for the union of 3 balls, one can calculate the volume of the union of 4 balls.
  2. 2.
    There is an extra point $x$ in $\Phi$ that follows $x_{3}$ in $\Phi$ and such that $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$. The case where $d(x,x_{2},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$ is symmetric.
    For evaluating the frequency of this event, we have to enrich $\cal D$ by adding $x$ and state that $x$ follows $x_{3}$ and that none of the points $x_{1},x_{2},x_{3},x_{4}$ follow $x$.
    $\begin{split}d(x,x_{3})<d(x,x_{1}),\,d(x,x_{3})<d(x,x_{2}),\,d(x,x_{3})<d(x,x_{4}),\\ d(x_{1},x_{3})<d(x_{1},x),\,d(x_{2},x_{3})<d(x_{2},x),\,d(x_{4},x_{3})<d(x_{4},x),d(x_{3},x_{4})<d(x_{3},x).\end{split}$ (15)
    This adds 7 quadratic inequalities. Finally, the condition that $d(x,x_{1},\Phi_{1})<d(x_{1},x_{2},\Phi_{1})$, gives one more quadratic inequality:
    $d(x^{\prime},x_{1}^{\prime})<d(x^{\prime}_{1},x^{\prime}_{2}),$ (16)
    where $x^{\prime}=\frac{x+x_{3}}{2}$. So the frequency of this refinement can be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{2}$, which is a refinement of $\cal D$} with one more variable and 8 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{5}$, the mean measure of $N_{1,2}^{(5)}$, which is the point process of the 5-tuples of points satisfying the conditions above, is given by,
    $\mathbb{E}[N_{1,2}^{(5)}(A)]$ $=$ $\mathbb{E}[\sum_{x,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}].$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,2}^{(5)}(A)]&=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1},\\ &1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}\Phi^{(5)}(dx\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    By the higher order Campbell-Little-Mecke formula,
    $\begin{split}\mathbb{E}[N_{1,2}^{(5)}(A)]&=\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{2}}\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}[1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}\\ &1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow x_{3}))=1}]\lambda^{(5)}(dxdx_{1}dx_{2}dx_{3}dx_{4}),\end{split}$
    where $\mathbb{E}^{x,x_{1},x_{2},x_{3},x_{4}}$ is the Palm expectation of $\Phi$ at $x,x_{1},...x_{4}$, and $\lambda^{(5)}$ is the factorial Poisson moment measure of order 5. Following the analogous steps as in case one, stationarity and the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,2}^{(5)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{2}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow 0)))}\lambda^{5}d\tilde{x}\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with $\tilde{{\cal D}_{2}}=\{(x-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},\,0,\,x_{4}-x_{3}):(x,\,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{2}\}$.
    Again, one can calculate the volume of the union of 4 balls.
  3. 3.
    There are two extra points $x$ and $y$, with $x\neq y$, such that $x$ follows $y$ (4 inequalities) and none of the points $x_{1},\ldots,x_{3}$ follow either $x$ or $y$ (6 inequalities). Thus we have
    $d(x,y)<d(x,x_{1}),\,d(x,y)<d(x,x_{2}),\,d(x,y)<d(x,x_{3}),\,d(x,y)<d(x,x_{4}),$
    $d(x_{1},x_{3})<d(x_{1},x),\,d(x_{1},x_{3})<d(x_{1},y),$
    $d(x_{2},x_{3})<d(x_{2},x),\,d(x_{2},x_{3})<d(x_{2},y),$
    $d(x_{3},x_{4})<d(x_{3},x),\,d(x_{3},x_{4})<d(x_{3},y).$
    To this, as above one should add the condition that $x^{\prime}$ is closer from $x_{1}^{\prime}$ than $x_{1}^{\prime}$ to $x_{2}^{\prime}$. WLOG take $x^{\prime}=\frac{x+y}{2}$ to be closer to $x^{\prime}_{1}$ than $x^{\prime}_{1}$ to $x^{\prime}_{2}$, i.e.,
    $d(x^{\prime}_{1},x^{\prime})<d(x^{\prime}_{1},x^{\prime}_{2}).$
    This amounts to 2 more variables and 11 more quadratic inequalities and one more empty ball conditions ($\Phi(B(x\to y))=1$). So the frequency of this refinement can be reduced to the evaluation of an integral over a semi-algebraic set ${\cal D}_{3}$, which is a refinement of $\cal D$} with two more variables and 11 more quadratic inequalities, with the function to be integrated involving one more ball in the union.
    For $A\in\mathcal{B}^{6}$, the mean measure of $N_{1,3}^{(6)}$, which is the point process of the 6-tuples of points satisfying the conditions above, is given by
    $\mathbb{E}[N_{1,3}^{(6)}(A)]$ $=$ $\mathbb{E}[\sum_{x,y,x_{1},x_{2},x_{3},x_{4}\in\Phi}^{\neq}1_{x,y,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{3}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1}$
    $1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}1_{\Phi(B(x\rightarrow y))=1}],$
    In integral form, this is
    $\begin{split}\mathbb{E}[N_{1,3}^{(6)}(A)]&=\mathbb{E}[\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}1_{x,y,x_{1},x_{2},x_{3},x_{4}\in A\cap{\cal D}_{3}}1_{\Phi(B(x_{1}\rightarrow x_{3}))=1},1_{\Phi(B(x_{2}\rightarrow x_{3}))=1}1_{\Phi(B(x_{3}\rightarrow x_{4}))=1}\\ &1_{\Phi(B(x\rightarrow y))=1}\Phi^{(6)}(dx\times dy\times dx_{1}\times dx_{2}\times dx_{3}\times dx_{4})].\end{split}$
    Analogously to case one, the calculation boils down to having $A=\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2}$ with $C$ a compact. Because we have a stationary point process, using the change of variables $\tilde{x}_{1}=x_{1}-x_{3}$, $\tilde{x}=x-x_{3}$, $\tilde{y}=y-x_{3}$ $\tilde{x}_{2}=x_{2}-x_{3}$, and $\tilde{x}_{4}=x_{4}-x_{3}$, we get
    $\mathbb{E}[N_{1,3}^{(6)}(\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\times C\times\mathbb{R}^{2})]$ $=|C|\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}\int_{{\mathbb{R}^{2}}}1_{\tilde{x},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4}\in\tilde{{\cal D}_{3}}}$
    $e^{-\lambda(\mathrm{Vol}(B(0\rightarrow\tilde{x}_{4})\cup B(\tilde{x}_{1}\rightarrow 0)\cup B(\tilde{x}_{2}\rightarrow 0)\cup B(\tilde{x}\rightarrow\tilde{y})))}\lambda^{6}d\tilde{x}d\tilde{y}d\tilde{x}_{1}d\tilde{x}_{2}d\tilde{x}_{4},$
    with
    $\tilde{{\cal D}_{3}}=\{(x-x_{3},\,y-x_{3},\,x_{1}-x_{3},\,x_{2}-x_{3},0,x_{4}-x_{3}):(x,y,x_{1},\,x_{2},\,x_{3},\,x_{4})\in{\cal D}_{3}\}.$
Let $\beta_{i}$ the frequency of event $i=1,2,3$ in the last list. Then the frequency of interest is
$\frac{1}{2}(\beta^{1}-\sum_{i=1}^{3}\beta_{i}).$ (17)
Note that the coefficient $\frac{1}{2}$ comes from a symmetry argument and to avoid double counting. We use numerical methods to evaluate all the frequencies. This is discussed in the next section.
5 Numerical Estimation Methods
We use two different methods to estimate frequencies. The first is discrete event spatial simulation, which consist in sampling a Poisson point process and exploiting the ergodic properties of the Poisson point process and related factors of such point processes [49, 7]. This approach is grounded in the ergodic theory of point processes [49]. The second method involves the numerical evaluation of integrals based on semi-algebraic sets discussed above. We explain both methods in detail and provide estimated values for some of the frequencies mentioned earlier. For each, we first provide some theoretical background and then describe how we apply it for our evaluation. When appropriate, we include 95% confidence intervals. Finally, we compare the results obtained from both methods.
5.1 Discrete Event Spatial Simulation
The background about ergodicity discussed in this section comes from [7].
Definition 5.
[7] A stationary framework $(\Omega,\mathcal{A},\{\theta_{t}\}_{t\in\mathbb{R}^{d}},\mathbb{P})$ is said to be ergodic if
$\lim_{a\rightarrow\infty}\frac{1}{(2a)^{d}}\int_{[-a,a]^{d}}\mathbb{P}(A_{1}\cap\theta_{-x}A_{2})dx=\mathbb{P}(A_{1})\mathbb{P}(A_{2}),\,\,\forall A_{1},A_{2},\in\mathcal{A}.$
It is said to be mixing if
$\lim_{|x|\rightarrow\infty}\mathbb{P}(A_{1}\cap\theta_{-x}A_{2})=\mathbb{P}(A_{1})\mathbb{P}(A_{2}),\,\,\forall A_{1},A_{2}\in\mathcal{A}.$
Lemma 1.
For all $n$, the follower point process of order $n$ is a factor of a Poisson point process, and is hence ergodic and mixing.
Discrete event spatial simulation (DESS) starts with sampling a Poisson point process in the fixed bounded window $W\subset\mathbb{R}^{2}$ (the observation window). We pick a convex set $W$, typically a square. In order to simulate a homogeneous Poisson point process, we use the ideas from [49]. In [49] it is suggested that simulating in a compact region $W$ can be split into two parts. Namely, the number of points in $W$ is determined from the simulation of a Poisson random variable. Then the positions of points in $W$ are obtained from simulating a binomial point process with that number of points. Since we are interested in the point process in the whole of $\mathbb{R}^{2}$, we take $W$ big. Hence there are small boundary effects.
Then we iteratively apply the Follower Dynamics to all the points simultaneously. We still have the mixing property satisfied. The first set of simulations is done using MATLAB and applying the Monte Carlo method. For each agent, we keep track of its leader at every step. In order to calculate any frequency, we count the number of agents in the window for which our sets of conditions are satisfied, and we divide it by the total number of agents in the window.
For example, for the frequency of leader pairs of order 0, we check for all pairs of agents and count the number of mutually closest neighbors at step 0. Since we keep track of the leader for each agent, we just check whether two agents are each others’ leaders.
Similarly, in order to determine the frequency of leader pairs of order 1, type 1, we check, for all pairs of agents, whether they had the same leader (but were not each others’ leaders) at step 0 and are each others’ leaders at step 1.
For the frequency of leader pairs of order 1, type 2, we check, for all pairs of agents, whether one was a leader of the other at step 0 and both are each others’ leaders at step 1.
Table 1 contains frequencies and $95\%$ confidence interval obtained using discrete event spatial simulation. In order to obtain a $95\%$ confidence interval we run simulation many times and keep track of the statistics. The statistics is obtained from $40$ samples containing in mean 20000 points each. For instance, observe that the frequency of leader pairs of order 1 is in the interval $[0.021,0.023]$ with $95\%$ confidence, which is the sum of both types shown in Table 1. Similarly, the frequency of inversion is in the interval $[0.13,0.14]$ with the same confidence.
Frequency of $95\%$ Confidence interval
Leader pair of order 0 $[0.6203,0.6227]$
Order 1 type 1 $[0.005,0.006]$
Order 1 type 2 $[0.0100,0.0105]$
4 body swap $[0.000063,0.0001]$
Inversion Type 1 $[0.138,0.140]$
Inversion Type 2 $[0.00019,0.00027]$
Table 1: Frequencies and $95\%$ confidence intervals obtained using DESS
5.2 Integrals on Semi-Algebraic Sets
5.2.1 Union of Balls
When exploiting our integral geometry representations of frequencies in the examples explored in Section 4, we end up with expressions that are integrals involving unions of balls. Hence, this way to find the probabilities of interest requires to compute the volumes of unions of balls. In this subsection, we first represent these unions of balls as semi-algebraic sets.
The setting is $\mathbb{R}^{d}$. Let $k$ be a positive integer. Let $c_{1},\ldots,c_{k}$ be arbitrary points of $\mathbb{R}^{d}$ which are the centers of balls of positive radii $r_{1},\ldots,r_{k}$, respectively. We are interested in the volume of
$\mathbb{U}:=\cup_{i=1}^{k}B(c_{i},r_{i}),$
with $B(c,r)$ the closed ball of center $c$ and radius $r$.
Let $\mathcal{S}$ be set of all non-empty subsets of $1,\ldots,k]$. For all $s\in S$, let
$\mathbb{V}_{s}=\cap_{i\in s}^{k}B(c_{i},r_{i})\cap_{j\notin s}^{k}B(c_{j},r_{j})^{c}.$
We have $\mathbb{U}=\cup_{s\in S}\mathbb{V}_{s},$ where the last union is disjoint.
Hence
$\mid\mathbb{U}\mid=\sum_{s\in S}\mid\mathbb{V}_{s}\mid.$
But for all $s$, $\mathbb{V}_{s}$ is the following semi-algebraic set:
[…]
Hence $\mathbb{U}$ is the disjoint union of the $2^{k}-1$ semi-algebraic sets $\{\mathbb{V}_{s},s\in S\}$ and its volume is the sum of the volumes of these semi-algebraic sets of $\mathbb{R}^{d}$.
5.2.2 Numerical evaluation of $\beta^{1}$
In this subsection, we derive the integral representation for the quantity $\beta^{1}$ using the semi-algebraic characterization of the underlying geometric constraints. The resulting six-dimensional integral is then evaluated numerically by Monte Carlo quadrature.
Semi-algebraic characterization of the integration domain
By stationarity, we fix $x_{3}=0$, obtaining
$\beta^{1}=\lambda^{4}\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathbf{1}_{(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4})\in\mathcal{D}_{0}}\,e^{-\lambda|\mathbb{U}(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4})|}\,d\tilde{x}_{1}\,d\tilde{x}_{2}\,d\tilde{x}_{4},$ (18)
where $\mathcal{D}_{0}=\mathcal{D}\cap\{x_{3}=0\}.$ Writing $r_{i}=|\tilde{x}_{i}|,d_{ij}=|\tilde{x}_{i}-\tilde{x}_{j}|,$ the defining conditions become
$r_{1}<d_{12},\qquad r_{1}<d_{14},\qquad r_{2}<d_{12},\qquad r_{2}<d_{24},$ (19)
$r_{4}<r_{1},\qquad r_{4}<r_{2},\qquad d_{12}<d_{14},\qquad d_{12}<d_{24}.$ (20)
The volume of the geometric exclusion region is
$|\mathbb{U}(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4})|=\mathrm{Vol}\!\Big(B(0,r_{4})\cup B(\tilde{x}_{1},r_{1})\cup B(\tilde{x}_{2},r_{2})\Big),$ (21)
namely the area of the union of the three disks associated with the nearest-neighbor constraints.
Writing
$D_{0}=B(0,r_{4}),\qquad D_{1}=B(\tilde{x}_{1},r_{1}),\qquad D_{2}=B(\tilde{x}_{2},r_{2}),$
the inclusion–exclusion formula gives
$|\mathbb{U}|$ $=|D_{0}|+|D_{1}|+|D_{2}|-|D_{0}\cap D_{1}|-|D_{0}\cap D_{2}|-|D_{1}\cap D_{2}|+|D_{0}\cap D_{1}\cap D_{2}|.$ (22)
After translating $x_{3}$ to the origin, $d(0,\tilde{x}_{1})=r_{1},d(0,\tilde{x}_{2})=r_{2},$ so all pairwise overlaps depend only on $r_{1},r_{2},r_{4},d_{12}$. Pairwise intersection areas are evaluated using the exact lens formula for disk intersections [21].
The integration domain $\mathcal{D}_{0}$ also admits an exact semi-algebraic description. Indeed, after squaring the inequalities in (20), each constraint becomes polynomial in the coordinates of $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{4})\in\mathbb{R}^{6}$. Thus $\mathcal{D}_{0}$ is a semi-algebraic subset of $\mathbb{R}^{6}$ in the sense of real algebraic geometry.
The numerical procedure therefore separates naturally into two parts:
  1. 1.
    an exact geometric characterization of the region satisfying the nearest-neighbor constraints through polynomial inequalities and exact overlap formulas;
  2. 2.
    a numerical approximation of the resulting integral by Monte Carlo quadrature.
Note that Monte Carlo sampling is used only for numerical integration and not for identifying the region itself. Belonging to $\mathcal{D}_{0}$ is checked exactly for each sampled configuration by evaluating the inequalities in (20).
5.2.3 Monte Carlo evaluation of the integral
The integral (18) is six-dimensional and involves nonlinear geometric constraints. In this setting Monte Carlo quadrature is considerably more effective than deterministic quadrature schemes.
The integration domain is described explicitly by the polynomial inequalities (20), so the semi-algebraic structure is used to determine exactly whether a sampled configuration belongs to $\mathcal{D}_{0}$. The Monte Carlo procedure is used only to approximate the resulting integral numerically.
We first rewrite the integral in polar coordinates. For each point
$\tilde{x}_{i}=(x_{i},y_{i})\in\mathbb{R}^{2},$
we introduce the change of variables
$\tilde{x}_{i}=(r_{i}\cos\theta_{i},r_{i}\sin\theta_{i}),\qquad r_{i}\geq 0,\quad\theta_{i}\in[0,2\pi).$
The Jacobian determinant of this transformation is $r_{i}$, so
$d\tilde{x}_{i}=r_{i}\,dr_{i}\,d\theta_{i}.$
The transformed integral is evaluated by Monte Carlo quadrature over the truncated parameter domain
$\mathcal{Q}=[0,R_{\max}]^{3}\times[0,2\pi)^{3},$
with $R_{\max}=7$. The truncation error decays exponentially because
$|\mathbb{U}|\geq\pi\max(r_{1}^{2},r_{2}^{2},r_{4}^{2}),$
so the integrand decays at least as fast as
$e^{-\lambda\pi r^{2}}.$
Sample configurations
$\xi^{(k)}=(\tilde{x}_{1}^{(k)},\tilde{x}_{2}^{(k)},\tilde{x}_{4}^{(k)})$
are generated independently in polar coordinates. For each point, the angular variable $\theta_{i}$ is sampled uniformly on $0,2\pi)$. The radial variables are sampled from the density
$f_{R}(r)=\frac{2r}{R_{\max}^{2}},\qquad 0\leq r\leq R_{\max},$
which is exactly the radial distribution induced by uniform area measure on the disk $B(0,R_{\max})$ expressed in polar coordinates. This is exactly the radial distribution induced by uniform sampling on the disk $B(0,R_{\max})$ expressed in polar coordinates. For each sampled configuration satisfying (20), we compute
$e^{-\lambda|\mathbb{U}(\xi^{(k)})|}.$
The estimator is
$\hat{\beta}^{1}=\lambda^{4}\frac{1}{A}\sum_{k=1}^{A}\mathbf{1}_{\xi^{(k)}\in\mathcal{D}_{0}}\,w(\xi^{(k)})\,e^{-\lambda|\mathbb{U}(\xi^{(k)})|},$
where $A=30000$ denotes the number of sampled configurations in each batch and $w(\xi^{(k)})$ is the corresponding density-correction factor arising from the polar sampling distribution.
To estimate variability, we perform $n=10$ independent batches.
The procedure is rigorous because (18) is an ordinary Lebesgue integral over $\mathbb{R}^{6}$ with explicitly specified indicator constraints. The estimator above is the standard Monte Carlo approximation of this integral by empirical averaging. Since
$0\leq\mathbf{1}_{\xi\in\mathcal{D}_{0}}e^{-\lambda|\mathbb{U}(\xi)|}\leq 1,$
the integrand is bounded and integrable. Consequently, the strong law of large numbers implies that
$\hat{\beta}^{1}\to\beta^{1}\qquad\text{almost surely as }A\to\infty.$
5.2.4 Comparison
As an independent validation of the integral computation, we estimate the same quantity using the discrete event spatial simulation (DESS) described in Section 5.1.
Numerical results.
Table 2 reports the resulting estimates for $\lambda=1$. The close agreement between the integral estimator and the DESS provides a consistency check for both the geometric characterization of the configuration space and the numerical implementation of the two procedures.
Table 2: Comparison of two independent estimators of $\beta^{1}$ for $\lambda=1$.
Method Estimate 95% confidence interval
Integral 0.00036 [0.00031, 0.00041]
DESS 0.00035 [0.00031, 0.00040]
6 Asymptotic Regime
This section focuses on the limiting behavior of parties. We analyze stable parties in Subsection 6.1. Stable parties are special types of Parties, which are trees for which the graph structure doesn’t change with steps. Such parties have a special structure. As we shall see they have no branching outside the root. We conjecture that all the Follower Parties eventually become stable parties.
Finally, in Subsection 6.2, we present some general long-term relations. As the number of steps in the Follower Dynamics approaches infinity, two sub-processes of the initial point process emerge. One converges in total variation - it is the ultimate leader pair point process, while the other converges weakly to its limit, the ultimate follower point process. Simulations in two dimensions show that the density of ultimate leader pairs is approximately 0.66. The remaining points, the ultimate followers, converge weakly to one of the ultimate leader pairs.
6.1 Stable Party
In this Section we focus on stable parties like in Figure 9, and show that for such parties we can derive exact positions of agents at any step. A simple example of a stable party is one leader pair and one follower.
Figure 9: Left:Tree at the initial time step. Right: Zoomed in situation after 20 time steps
After one step of the dynamics, the leader pair becomes fixed. We therefore index all positions relative to this step. Let $a_{1}=a_{1}^{(1)}$ denote the (fixed) leader, and let $a_{n}=a_{n}^{(1)}$ denote the position of its follower of order n1n-1. The position of agent $n$ at time $i$ is denoted by $a_{n}^{(i)}$.
The dynamics are governed by the same rule of moving halfway to the leader’s position
$a_{n}^{(i)}=\frac{1}{2}\big(a_{n}^{(i-1)}+a_{n-1}^{(i-1)}\big),$
so each position remains a weighted sum of the initial data.
Lemma 2.
If the party is stable, then for every $n\geq 1$ and $i\geq 1$,
$a_{n}^{(i)}=\begin{cases}\frac{1}{2^{i}}\displaystyle\sum_{k=0}^{i}{i\choose k}a_{n-k},&i<n,\\[10.00002pt] \frac{1}{2^{i}}\left(\displaystyle\sum_{k=0}^{n-1}{i\choose k}a_{n-k}+\sum_{k=n}^{i}{i\choose k}a_{1}\right),&i\geqslant n.\end{cases}$
Proof.
The claim follows by induction on $i$, using the linearity of the update rule. ∎
In particular, for fixed $n$,
$\frac{1}{2^{i}}\sum_{k=0}^{n-1}{i\choose k}a_{n-k}\to 0,\qquad\frac{1}{2^{i}}\sum_{k=n}^{i}{i\choose k}a_{1}\to a_{1},\quad\text{as }i\to\infty,$
so every agent converges to the leader.
Proposition 1.
If the party is stable, then it cannot have branching away from the root.
Proof.
Suppose that branching occurs at some node $a_{n}$ with $n\geq 2$, and let $a_{n+1}^{(1)}$ and $a_{n+1}^{(2)}$ denote its two children.
By linearity, common ancestry terms cancel, and the distance between siblings evolves as
$d\big(a_{n+1}^{(1)}(i),a_{n+1}^{(2)}(i)\big)=\frac{1}{2^{i}}d\big(a_{n+1}^{(1)},a_{n+1}^{(2)}\big).$
On the other hand, using Lemma 2,
$a_{n+1}^{(1)}(i)-a_{n}^{(i)}=\frac{1}{2^{i}}\sum_{k=0}^{n-1}\binom{i}{k}\big(a_{n+1-k}^{(1)}-a_{n-k}\big).$
The leading term corresponds to k=n1k=n-1 and grows like $\binom{i}{n-1}$, while all remaining terms are of strictly lower order. Hence
$\big|a_{n+1}^{(1)}(i)-a_{n}^{(i)}\big|=\frac{1}{2^{i}}\big(c\,i^{n-1}+O(i^{n-2})\big),$
for some $c>0$.
Comparing the two distances, both decay at rate $2^{-i}$, but the distance to the parent carries an additional polynomial factor $i^{n-1}$. Therefore, for all sufficiently large $i$,
$d\big(a_{n+1}^{(1)}(i),a_{n}^{(i)}\big)>d\big(a_{n+1}^{(1)}(i),a_{n+1}^{(2)}(i)\big),$
contradicting stability. ∎
Remark. Linearity implies that sibling distances contract purely exponentially, while distances to ancestors retain polynomial memory of the initial configuration, forcing eventual separation.
The proposition suggests that stable configurations are necessarily chain-like. This leads to the following conjectures.
Conjecture 1.
For every leader pair of any order, there exists a time after which the associated party becomes stable.
Conjecture 2.
Starting from a Poisson point process in the plane, the limiting configuration consists only of ultimate leader pairs and their followers arranged in chains.
6.2 Long Term Relations
In this subsection we look at some mass transport relations that hold in general in the limiting process. Limiting objects consist of ultimate leader pairs and ultimate followers, we can derive a relation between them. Let $\lambda=1$ be the intensity of the initial Poisson point process. Let $\lambda_{f}$ be the intensity of the ultimate followers and $\lambda_{l}$ be the intensity of the ultimate leaders. Take $\bar{N}$ to be the mean number of ultimate followers of one ultimate leader. Then by the mass transport principle it directly follows that
$1\cdot\lambda_{f}=\bar{N}\lambda_{l},$
using that $\lambda_{f}+\lambda_{l}=1$, implies that
$\bar{N}=\frac{1}{\lambda_{l}}-1.$
Since there are two ultimate leaders in a pair, the total number of ultimate followers is $N=2\bar{N}$. This relation should hold regardless of the dimensions.
Now if we plug in $\lambda_{l}=\frac{2}{3}$, like obtained in the simulations, we get that $N=1$. Implying that the mean size of a party is 3 in the limit, which is also consistent with the simulations.
Conjecture 3.
There exist ultimate followers.
7 Conclusion
In summary, we introduced a new model inspired by problems in opinion dynamics. We described various phenomena related to this dynamics and examined the system’s long-term behavior. We provided a systematic way to calculate the frequencies of specific configurations, such as the densities of certain leader pairs. Additionally, we numerically analyzed the long-term behavior of parties and derived general results about the shapes of limiting parties. For future work, we plan to prove that parties remain finite at all steps, supported by simulations. Furthermore, we aim to demonstrate that the limiting objects can only be either ultimate leader pairs or ultimate followers. We also plan to explain the observed ratio between the densities of the ultimate leader pair point process and the ultimate follower point process. Finally, we plan to extend these results to higher dimensions.